### v2

# A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies

Abstract
Understanding how social norms vary across cultures can help us build culturally aligned NLP systems. We propose a culture-agnostic approach to norm discovery, using moral emotions, shame and pride, to identify examples of normative expectations and extract corresponding social norms. We present the first cross-cultural self-conscious emotions dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over 10K extracted social norms. We validate our dataset using native speakers and demonstrate how our dataset reveals variations in social norms that align with the cultural dichotomy observed in these nations -- e.g., Bollywood movies emphasize shame due to deviation from social roles, and express pride in family honor, while Hollywood shames poverty and incompetence, and takes pride in ethical behavior. Notably, females are shamed more across both cultures and both cultures shame women for violating similar normative expectations.
A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies
1 Introduction
Normative expectations, [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al. (1990) use injunctive norms. We use the terms social norms and normative expectations interchangeably. commonly called ‘‘social norms’’, refer to people’s beliefs about what others approve of or expect people to do. These norms play a significant role in shaping and regulating social behavior and can reveal underlying moral and ethical values.
Recent research on social alignment of LLMs has either crowd-sourced social norms from QnA style posts on social media platforms or prompted pre-trained LLMs to extract them Jiang et al. (2021); Nahian et al. (2020); Fung et al. (2022). However, a majority of these studies are conducted in English and capture Western social situations. Furthermore, LLMs such as GPT-4 OpenAI (2023) are predominantly trained on data from Western cultures, Havaldar et al. (2023b), and thus, the generated "social norms" tend to be Anglocentric.
Look, Alok is in love with me and I love him too.
Shameless one! Have you lost all your shame?
What is there to be ashamed in this? Don’t sisters love their brothers in this vicinity?
Okay, I am going.
$\Rightarrow$ Incestuous relationship evokes shame
Sister-in-law! Congrats, sister-in-law! Big brother has started working!
Really?
Now you will have a place of pride in this family!
Yes, please!
Transfer 10 lakhs, okay? Bye!
$\Rightarrow$ Employed husband evokes pride
Table 1: Excerpts of dialogues with target search words.
Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
Self-conscious emotions, or moral emotions, are emotional reactions to a violation or an observance of normative expectations in society. In turn, these emotions reflect social sanctions and approvals. Unlike basic emotions such as joy and anger, self-conscious emotions such as shame and pride are evoked by self-reflection and self-evaluation. Examining situations evoking self-conscious emotions can thus reveal unspoken social norms specific to a culture (See Table 1). While the function (i.e., social regulation) of self-conscious emotions is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022), their appraisals and expressions vary Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022).
In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, (b) propose a culture-agnostic approach for norm discovery and release a dataset of $10K$ normative expectations and, (c) examine cross-cultural variations in social norms. We study India and the United States of America to investigate the utility of moral emotions for cross-culture norm discovery. India is predominantly a collectivist society where one’s sense of self is interwoven with community beliefs, whereas the USA is predominantly an individualist society that values competency and autonomy Triandis (1989, 1988). The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their moral and ethical values.
Relatedly, gender differences in social behaviors emerge from societal expectations and the process of socialization. For instance, women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We examine how shame, an indicator of negative evaluation, is attributed to women across cultures. We hypothesize that more shame is attributed to women in Indian society due to its patriarchal nature compared to American society.
Movies Total Tokens #shame #pride #control
Bollywood 2738 22.62M 5409 2999 8303
Hollywood 2697 20.78M 1221 2805 4385
5435 43.3M 6583 5804 12688
Table 2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
2 Background
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. SOCIAL-CHEM-101 Forbes et al. (2020), SCRUPLES Lourie et al. (2021), SOCIAL BIAS FRAMES Sap et al. (2019), MORAL INTEGRITY CORPUS Ziems et al. (2022), and VALUE PRISM Sorensen et al. (2023) are a few datasets developed to teach socially aligned interactions to LLMs. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations in these datasets and the human annotators employed to label those situations reflect English beliefs and ethics. Additionally, language models prompted to identify "social norms" are known to reflect the values and beliefs of WEIRD nations (Western, Educated, Industrialized, Rich, and Democratic) Havaldar et al. (2023b); Atari et al. (2023). Capturing diversity in morality is thus a research question of great significance Talat et al. (2021).
In this paper, we examine the cross-cultural variations in self-conscious emotions and use these emotions to extract culture-specific normative expectations. We validate our findings by evaluating their alignment with known cultural tendencies of Indian and American societies. Additionally, we examine gender-wise attribution of shame and norm distribution to contextualize the potential impact of inculcating automatically extracted social values into machines.
3 Dataset
Culture-specific people’s conversations are needed to study moral emotions and to mine underlying normative expectations. Social media QnA posts are primarily recollections or ’what-if’ situations that do not reveal fully how these emotions are expressed in natural conversations, for instance between a boss and an employee or among family members at a Christmas table.
Movies are a rich source of real-time social interactions illustrating power dynamics, negotiations, and self-conscious emotions in action; to study how and why these emotions are expressed. We collected English subtitles for 5,435 movies belonging to Bollywood and Hollywood that were released post-1990 by automatically crawling websites that host or link movie subtitles (See Table 2 for movie distribution). The choice of English language for Bollywood movies was made based on the availability of subtitles. The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
3.1 Data Proprocessing
We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context. One author manually checked the entire dataset of shame and pride-related dialogues to correct spelling mistakes in subtitles, remove unrecognizable characters, and identify phrases having pride and shame-related keywords but not indicating social sanction (or approval).
For instance, the phrases like "what a shame" or "it’s a shame" in Hollywood movies do not reflect social disapproval, and therefore such dialogues were excluded. Below is an example of such dialogue.
‘‘…His music is sad all of the time. It’s really a shame. He’ll never be allowed to go up to heaven.   But you will, won’t you, Johan?   Oh yes. I’ll get there."
Similarly, the phrase "proudly presents" was used in the context of introducing an event or an artist in Hollywood movies. All dialogues with the phrase "proudly presents" were thus removed for further analysis.
We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Additionally, we also formed a control set of dialogues unrelated to shame and pride for both movie industries (See Table 2 for dialogue distribution).
Table 3: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with shame for Bollywood and Hollywood after Benjamini-Hochberg p-value correction. Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood. See Table A6 and A8 for the complete set of correlations.
4 Approach
Fig. 1 illustrates our study design comprising two key approaches (a) a vocabulary approach to measure cross-cultural variations in linguistic expressions of shame/pride and (b) prompting LLM approach to extract latent normative expectations and examine how they differ between India and the U.S.A.
4.1 Vocabulary Approach: LIWC
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool widely used in psychology to identify psychological properties of words such as "self-focus" (e.g., I, my, mine) vs. "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we extracted the normalized distribution of psycho-social categories from the dialogues (shame+control vs pride+control for both movie industries) and examined their correlation with shame and pride. Please note that search keywords (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
4.2 Pretrained LLM: GPT-4
Taking into account pre-trained LLMs’ tendency to favor an Anglocentric understanding of text Havaldar et al. (2023b), we designed the prompts to elicit the "reason behind self-conscious emotions" rather than extracting "social norm" to overcome the cultural bias. The prompts were designed to seek the following:
  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender?
  • What is the reason behind $<$the feeling of shame/pride$>$?
The first question orients the LLM to focus on the subject, i.e., person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social norm in the culture. Since GPT-4 exhibits performance at par with human annotators Pan et al. (2023); Törnberg (2023), we designed this task using GPT-4 chat in a two-shot setting. The temperature was set to 00. The prompts with two-shot samples are provided in Tables A4 and A5.
Themes in Social Norms
To understand differences in situations that evoke shame and pride in Indian and American societies, we performed agglomerative clustering after embedding extracted norms using SBERT embeddings Reimers and Gurevych (2019). For this purpose, we merged shame-related norms from Bollywood and Hollywood and pride-related norms from Bollywood and Hollywood and removed duplicates. Social norms extracted from GPT-4 chat are short phrases (See Table 5) and are devoid of culture-specific style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
Table 4: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction. See Tables A7 and A9 for all significantly correlated categories.
5 Results
5.1 Cross-cultural variations in expressions of shame and pride
Shame
The selected top LIWC categories significantly correlated with shame in Bollywood and Hollywood is provided in Table 3. Although shame is associated with negative emotions, power, and morality in both industries (See Table A6 and A8 for all significant correlations), significant cultural nuances exist in its manifestation. In Hollywood, shame is I-focused with remorse (e.g., sadness, negative emotions, anxiety), whereas in Bollywood, it is You-focused with more social references (e.g., you, your, he, her). Shame is past-focused in Hollywood, whereas it is present-focused in Bollywood movies. Anxiety is unique to Hollywood-shame whereas anger is more strongly associated with shame in Bollywood. Categories such as female, sexual, and swearing are exclusively correlated with shame in Bollywood indicating a conservative society whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
Pride-
related interactions are family focused and self-centered (e.g., I am proud of __) in Hollywood movies whereas it is achievement/power focused and "We"-centered in Bollywood movies (See Table 4). Similarly, pride expressions related to power, morals, and prosocial behavior are more common in Bollywood dialogues, whereas pride expressions related to individuals’ growth (i.e., male/female, work) are more frequent in Hollywood dialogues. Both industries associate pride with ethnicity.
5.2 Cross-cultural variations in Social norms
Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing the self-conscious emotions (shame and pride)" and the reason behind self-conscious emotions in a randomly sampled set of $100$ dialogues from Hollywood and $100$ dialogues from Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
Gender was incorrect for eight samples in the Bollywood set and for five samples in the Hollywood set. For Bollywood, GPT-4 chat predicted a gender for 15 (3 predicted as Female, rest as male) samples where gender was not evident, whereas, for 10 (3 Female) such samples in Hollywood. One such example is provided below:
e.g. We’re proud of Bristol’s decision to have her baby… and even prouder to become grandparents? We’re not proud that… our teenage daughter is pregnant, Maria. And I don’t want to send a message that teen pregnancy… is something to be proud of. I want that line out. Of course, Governor.
Here, the gender is predicted as "Male" for Governor. GPT-4 chat returns the most likely gender indicating bias however, it is not always incorrect. It is thus safe to assume that prediction error for gender lies between 8-23% for Bollywood and 5-15% for Hollywood.
For Bollywood, 11 predicted reasons behind self-conscious emotion were incorrect whereas for Hollywood, 2 such samples were incorrect. The reason was not explicit for another nine samples in Bollywood set and one sample in Hollywood. Overall, the prediction error lies between 11-20% for Bollywood whereas it is 2-3% for Hollywood.
Bollywood
eavesdropping on private conversation
expressing love for a man
Incestuous relationship
giving birth to a girl child
Poor academic performance
Hollywood
not living up to expectations
hiding/avoiding confrontation
not returning calls after intimacy
mistreatment of a woman
offering poor quality goods for sale
Table 5: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
Dataset of cross-cultural social norms
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (99.2%) shame-related dialogues out of 5363, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is expressed even without a reason to express affection, specifically toward close family members such as son. GPT-4 chat predicted a total of 10,445 social norms (See Table A2 for norm distribution). A sample of five reasons extracted for shame and pride is provided in Tables 5 and 6.
Bollywood
mastering a trick
fulfilling father’s dreams
provide care for old parents
fiancee’s physical appearance
his wealth
Hollywood
for being a hard worker regardless of the task
being a brilliant student
winning olympic gold
achievements and growth
coming out as queer
Table 6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms.
5.2.1 Themes of Social Norms
We found twenty-five clusters for shame-related norms and twelve for pride-related norms using agglomerative clustering ( See Tables A12 and A13 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A10 and A11 for examples). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1.
$\vec{\Delta}=\forall_{t_{i}\in themes}\frac{\mathcal{D}_{bolly_{t_{i}}}}{{\mathcal{D}_{bolly}}}-\frac{\mathcal{D}_{holly_{t_{i}}}}{\mathcal{D}_{holly}}$ (1)
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{Holly,Bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis).
Shame-related social norms
Poverty and inappropriate behaviors are more strongly associated with shame in Hollywood whereas gender conformity and disrespect are more strongly associated with shame in Bollywood (See Fig. 2). As expected, collectivist factors such as parent-related roles and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as incompetence and poor social etiquette evoke shame in Hollywood.
Pride-related social norms
Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas serving the nation, family honor, and son’s success are associated with pride in Bollywood (See Fig. 3).
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
5.2.2 Gender differences in Attribution of Self-conscious Emotions
For the Bollywood set, GPT-4 predicted 1632 targets as female and 4361 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table A2), there are more male targets than females. This aligns with Geena Davis Inclusion Quotient (2016) findings that male characters get more screen time and speak more than female characters.
Using target gender as predicted by GPT-4 (See Table A2 for distribution), we computed the gender-wise attribution to emotions "shame" and "pride" using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
$\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is similar (0.16 for Hollywood and 0.18 for Bollywood) irrespective of the movie industry.
Normative expectations evoking shame are surprisingly similar for females across movie industries (See Figs. 5 and 6). Sexuality (e.g., promiscuity, immodesty, gender roles) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. For pride, women express more pride for family roles. However, we do see a variation in themes for men-pride i.e., justice, bravery and achievement are major themes for pride in Bollywood whereas it is duty, and self-identity in Hollywood. We also note interesting reverse patterns e.g., self-identity is a source of pride for females in Bollywood whereas it is for males in Hollywood.
Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
6 Discussion
Cross-cultural variations in expressions of moral emotions
The entailment of shame tells us about undesired behavior, whereas pride reveals desired behaviors in a society. While self-conscious emotions are culture-agnostic markers of unspoken social rules, the expressions of these emotions vary across cultures. We uncovered these variations in Sec 5.1. Shame is a highly undesirable self-focused emotion in the U.S. emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al. (2013). Notably, we also observe a strong correlation of first-person singular pronouns with shame and its infrequent use in Hollywood movies compared to Bollywood movies (See Table 2). In contrast, shame is interdependent in collectivist communities Wong and Tsai (2007). Our empirical analyses reveal You-focused (e.g. You, social references) nature of shame in Bollywood movies (See Table A6). Moreover, the contrasting tenses coupled with different emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). The correlation of shame with categories such as sexual, female, conflict, swear is exclusive to Bollywood, likely due to emphasis on conformity in Indian society. Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial).
Moral Emotion-based Prompting
Compared to GoEmotions dataset Demszky et al. (2020) having 817 samples for embarrassment and 452 samples for pride, we provide a much larger cross-cultural dataset of shame (6583 samples) and pride (5804 samples) related discourse, collected from movie subtitles across two cultures. Additionally, we release 10K culture-specific social norms. These can be used for downstream applications like evaluation of culturally aware language inference and achieving pluralistic values in LLMs.
Self-conscious emotion-based prompting helped in mitigating cultural bias in LLMs-generated text and potential stereotyping during norm discovery; as evident from the specificity of the norms extracted using our approach ( See Tables 5 and 6). More importantly, these norms reflect the cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter vs. honesty in business, returning calls after date night as in Table 5) revealing the differences in socially expected behaviors in Indian society and the U.S. society. The excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). A natural next step would be to capture situations with implicit shame and pride. The pipeline in Fig. 1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text. For this purpose, our dataset of shame/pride-related situations will be useful.
Cross-cultural variations in Social Norms
To tease apart underlying values evoking moral emotions across cultures, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012). However, a social norm can have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction. However, the diachronic shift in the underlying values of social norms is out of the scope of this paper. To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S. society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations. Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (i.e., nation honor, family honor) in Bollywood. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes still reflect the patterns seen during LIWC analysis, supporting the use of moral emotions for overcoming cultural bias in LLMs.
Gender Differences
Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies. Surprisingly, the gender gap in pride- and shame-related dialogues is similar across both societies. While our initial hypothesis that "more shame is attributed to women in Indian society compared to American society" (0.18 in Bollywood vs 0.16 in Hollywood) is true, the difference is negligible. It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies. Our analysis reveals that women are more frequent targets of social sanctions than men. It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
7 Conclusion
We introduce an emotion-based approach to extract and analyze social norms across cultures. We create the largest known multi-cultural dataset of self-conscious emotions and relevant underlying norms. Our post hoc analysis of social norms demonstrates (a) cross-cultural linguistic differences in self-conscious emotions (shame and pride), (b) cultural dichotomy in normative expectations in India and the U.S.A., and (c) more "social sanctions" and fewer "social benefits" to the female gender in movies.
Future work can utilize our dataset to better align LLMs to various cultures. Additionally, our analysis of self-conscious emotions is the first of its kind; we hope future NLP researchers will build upon this work to investigate social norms in LLMs from a multicultural, emotion-based perspective.
Social Impact and Ethics Statement
Social norms discovery is a crucial component in programs [^2] ALIGN-https://www.alignplatform.org/learning-collaborative designed for social and behavioral change to promote equity, social justice, and well-being Mauduy et al. (2022); Bonan et al. (2020). Further work explores style as a product of norm differences Havaldar et al. (2023a). Social psychology investigates social norms (descriptive vs injunctive) to design experiments for understanding behaviors such as self-regulation, persuasion Cialdini et al. (1990) and decision-making Gavrilets (2020); Bhanot (2021) to promote collective-level change in societies.
Relatedly, Kimbrough and Vostroknutov (2023) showed people’s tendency to choose self-serving social norms using a dictator-recipient setup, emphasizing the need for dedicated research efforts to understand morality and belief distortion in different contexts. The norms and cultural preferences learned from movies that often showcase stereotypical behaviors of society may induce pluralistic ignorance and, more importantly, lead to discrimination and biases in LLMs when used for training. We hope that this paper will encourage scrutiny of source corpora and derived norms before their use for fine-tuning LLMs.
Limitations
Social norms mutate as society evolves. We acknowledge that our dataset of movies (released post-1990) may reflect social norms that are less characteristic of contemporary society. Moreover, countries like India and America contain a mix of cultures. The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North India vs South India). Movies also exaggeratedly depict the world around us (e.g., weddings, criminal activities, sexual abuse, etc.), and we caution against stereotyping cultures based on movie-based norms.
The dominant language in Bollywood movies is Hindi and our analysis is based on their English translations which may not always be accurate, especially when the discourse is about concepts native to a culture. Relatedly, the LIWC may not have high coverage for such concepts. We did not compare the movie genre and acknowledge that situational/unrealistic norms (e.g., science fiction, comedy, etc.) could exist. Regardless, social norms associated with shame and pride are still relatable and reflect the target audience’s beliefs. It is also worth noting that gender norms in movies may not fully reflect the norms of today’s society. This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
The prompts ("You study Indian society" vs "You study Western society" described in Tables A4 and A5 ) used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
Appendix A Appendix
A.1 Search Keywords
Search keywords defined in Table A1 were used to extract shame- and pride-related dialogues from Bollywood and Hollywood movie subtitles.
Lemma Word forms
shame shame, shamed, shameful, ashamed
pride proud, prouder, proudly, pride
Table A1: List of search keywords for extracting dialogues.
A.2 Norms Distributions
Over 10k reasons (or norms) were extracted using GPT-4. The distribution of norms and gender of targets is provided in Table A2.
Bollywood Hollywood
Gender shame pride shame pride
male 3102 1259 591 776
female 1306 326 246 236
not known 913 652 319 719
total 5321 2237 1156 1731
Table A2: Norm and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female. The duplicate norms are not removed as their frequency reflect their prevalence and is useful for estimating gender association.
A.3 Annotation
The annotation guidelines to verify the gender predicted by GPT-4 and the correctness of the reason are provided in Table A3. The annotators for Bollywood set and Hollywood set were Indian and American respectively. Both annotators were female, proficient in English language, and well-versed with social norms. During annotation, if the gender or the reason is unclear, the annotators were asked to label "not explicit". We only considered the cases where the gender was predicted to be either male or female. The task is objective and inter-annotator agreement was not computed. The annotators volunteered for the task and were not provided monetary compensation.
Guidelines for Manual Evaluation
1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
2. Identify the gender. Check gender markers such as Mr/Mrs., s/he, him/her, etc. If the name is provided in the conversation, check if the name is likely to be a male name or female. If not clear, mark "not explicit".
3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
Table A3: Guidelines for Annotation
A.4 Prompts
The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A4 and A5.
Prompts for norm discovery in Bollywood movie dialogues
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
2. Identify the reason behind feeling proud. If not explicit, return NA.
The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
Input: Your mother had already given up on me because I was useless hopeless But you were the apple of her eye My child will make me proud look after me He’ll support in my old-age But this kid left her all alone.
Output: mother, male, provide care for old parents
Input: I’ve heard there’s a promising young student in your school What’s his name? He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre
Output: authority, NA, Sports achievement"
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
2. Identify the primary reason for shaming
The answer should be short and in CSV format. $<$ shamed, gender, reason$>$
Input: And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin.
Output: girl’s parent, NA, not able to marry off their daughters
Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’
Output: System, NA, poor law and regulations"
Table A4: Bollywood: Prompts for norm discovery using GPT-4 Chat.
Prompts for norm discovery in Hollywood movie dialogues
"You are an experienced social scientist and you study Western society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
2. Identify the reason behind feeling proud? If not explicit, return NA.
The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir?
Output: Sir, male, winning olympic gold
Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart.
Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study western societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
2. Identify the primary reason for shaming.
The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$
Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake.
Output: NA, male, not returning borrowed money
Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours.
Output: Snake, male, stealing and preying on people"
Table A5: Hollywood: Prompts for norm discovery using GPT-4 Chat.
A.5 LIWC Correlation Results
Tables A6, A7, A8 and A9 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
LIWC Categories Top-5 words correlation 95% CI
EMO_NEG (bad, mad, scared, worry, fear) 0.330 [0.315, 0.344]
TONE_NEG (lost, kill, wrong, bad, hit) 0.249 [0.233, 0.265]
POWER (sir, respect, own, kill, poor) 0.198 [0.182, 0.214]
EMOTION (love, good, bad, happy, crazy) 0.168 [0.152, 0.185]
YOU (you, your, you’re, yourself, you’ve) 0.161 [0.145, 0.178]
SOCREFS (you, your, he, her, him) 0.148 [0.132, 0.165]
FEELING (feel, touch, feeling, felt, hard) 0.133 [0.117, 0.150]
DRIVES (we, our, us, sir, married) 0.115 [0.098, 0.132]
SOCIAL (you, your, he, her, him) 0.110 [0.094, 0.127]
MORAL (wrong, innocent, duty, decent, excuse) 0.102 [0.085, 0.118]
AFFECT (love, good, keep, respect, well) 0.078 [0.062, 0.095]
NEGATE (not, don’t, no, aren’t, won’t) 0.065 [0.049, 0.082]
FEMALE (her, she, girl, she’s, mom) 0.060 [0.043, 0.077]
PPRON (you, i, me, your, my) 0.058 [0.041, 0.075]
FAMILY (son, married, uncle, dad, mom) 0.055 [0.039, 0.072]
PREP (to, of, in, for, on) 0.053 [0.037, 0.070]
SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 [0.034, 0.067]
PRONOUN (you, i, me, your, my) 0.040 [0.023, 0.057]
AUXVERB (is, are, have, be, don’t) 0.029 [0.013, 0.046]
CONFLICT (kill, killed, accusing, killing, cruel) 0.024 [0.007, 0.041]
SWEAR (hell, bloody, idiot, damn, ass) 0.024 [0.007, 0.041]
EMO_ANGER (mad, angry, hate, cruel, argue) 0.022 [0.005, 0.039]
FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 [0.005, 0.038]
Table A6: Psychosocial categories significantly correlated (p<0.05) with shame in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words correlation 95% CI
DRIVES (we, our, us, sir, work) 0.122 [0.104,0.140]
DET (the, a, my, your, that) 0.101 [0.082,0.119]
ACHIEVE (work, better, win, best, try) 0.098 [0.080,0.116]
POWER (sir, own, respect, kill, power) 0.091 [0.072,0.109]
SOCREFS (you, your, he, we, our) 0.085 [0.067,0.103]
PREP (of, to, in, for, with) 0.084 [0.066,0.103]
MORAL (wrong, duty, brave, arrogant, useless) 0.075 [0.057,0.093]
CONJ (and, but, so, if, as) 0.075 [0.056,0.093]
REWARD (win, won, glory, success, successful) 0.071 [0.052,0.089]
TONE _POS (love, good, thank, well, great) 0.064 [0.046,0.082]
POLITIC (nation, army, sultan, president, dynasty) 0.064 [0.045,0.082]
WE (we, our, us, we’ll, let’s) 0.060 [0.042,0.079]
SOCIAL (you, your, he, we, our) 0.060 [0.041,0.078]
FAMILY (son, papa, married, dad, uncle) 0.059 [0.041,0.078]
AFFILIATION (we, our, us, dear, we’ll) 0.059 [0.041,0.077]
FEELING (feel, feeling, hard, felt, sense) 0.055 [0.036,0.073]
ETHNICITY (indian, indians, british, hindi, caste) 0.054 [0.036,0.072]
MALE (he, his, him, son, sir) 0.054 [0.036,0.072]
CULTURE (indian, nation, army, car, indians) 0.044 [0.026,0.062]
AFFECT (love, good, thank, well, great) 0.040 [0.021,0.058]
ARTICLE (the, a, an, tha) 0.039 [0.021,0.058]
PPRON (you, i, my, your, me) 0.036 [0.017,0.054]
PROSOCIAL (thank, please, sorry, respect, gift) 0.032 [0.014,0.051]
FUNCTION (you, the, i, of, to) 0.032 [0.013,0.050]
YOU (you, your, you’re, you’ve, you’ll) 0.030 [0.011,0.048]
THEY (they, their, them, they’re, they’ll) 0.029 [0.011,0.048]
CERTITUDE (really, real, surely, proved, actually) 0.021 [0.002,0.039]
EMO _POS (love, good, happy, happiness, smile) 0.020 [0.001,0.038]
Table A7: Psychosocial categories significantly correlated (p<0.05) with pride in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words correlation 95% CI
EMO_NEG (sick, pain, fear, bad, afraid) 0.425 [0.403, 0.446]
TONE_NEG (lost, wrong, sick, pain, poor) 0.355 [0.331, 0.377]
EMOTION (good, love, sick, pain, bad) 0.290 [0.266, 0.314]
POWER (own, sir, poor, killed, war) 0.263 [0.238, 0.287]
AFFECT (well, good, love, help, damn) 0.168 [0.142, 0.193]
DRIVES (we, us, our, work, we’re) 0.152 [0.127, 0.178]
FUNCTION (you, i, the, to, of) 0.131 [0.105, 0.157]
PPRON (you, i, me, i’m, my) 0.111 [0.085, 0.137]
MORAL (wrong, excuse, decent, honest, duty) 0.110 [0.084, 0.136]
I (i, me, i’m, my, i’ll) 0.102 [0.076, 0.128]
EMO_SAD (crying, cry, sob, lonely, sad) 0.096 [0.070, 0.122]
NEGATE (no, not, don’t, nothing, never) 0.090 [0.063, 0.116]
PREP (to, of, in, for, on) 0.089 [0.062, 0.115]
PRONOUN (you, i, that, it, me) 0.088 [0.062, 0.114]
YOU (you, your, you’re, yourself, you’ve) 0.080 [0.054, 0.106]
AUXVERB (be, i’m, is, was, have) 0.077 [0.051, 0.103]
FOCUSPAST (was, did, were, been, didn’t) 0.069 [0.043, 0.096]
SOCIAL (you, your, we, he, you’re) 0.060 [0.033, 0.086]
CONJ (and, so, but, if, when) 0.056 [0.030, 0.082]
LINGUISTIC (you, i, the, to, of) 0.056 [0.030, 0.082]
ALLNONE (no, all, nothing, never, yes) 0.055 [0.029, 0.081]
FAMILY (son, dad, baby, mom, mama) 0.054 [0.028, 0.081]
SOCREFS (you, your, we, he, you’re) 0.050 [0.024, 0.077]
EMO_ANX (fear, afraid, worry, terrified, scared) 0.042 [0.015, 0.068]
ILLNESS (sick, pain, pains, flu, sickly) 0.038 [0.012, 0.064]
FEELING (feel, felt, pain, feeling, hard) 0.035 [0.008, 0.061]
EMO_ANGER (hate, hated, mad, angry, hates) 0.033 [0.007, 0.059]
DIFFER (not, but, if, didn’t, or) 0.033 [0.007, 0.059]
DISCREP (should, can, would, can’t, want) 0.032 [0.006, 0.059]
COGNITION (no, not, all, know, but) 0.032 [0.005, 0.058]
Table A8: Psychosocial categories significantly correlated (p<0.05) with shame in Hollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Words correlation 95% CI
PPRON (you, i, i’m, me, my) 0.135 [0.113,0.158]
SOCREFS (you, your, we, he, you’re) 0.131 [0.108,0.154]
FAMILY (son, dad, baby, mom, mama) 0.121 [0.099,0.144]
CONJ (and, so, but, as, if) 0.114 [0.091,0.137]
SOCIAL (you, your, we, he, you’re) 0.102 [0.079,0.125]
FUNCTION (you, i, the, of, to) 0.100 [0.077,0.123]
I (i, i’m, me, my, i’ll) 0.095 [0.072,0.118]
YOU (you, your, you’re, you’ve, yourself) 0.091 [0.068,0.114]
MALE (he, his, him, man, son) 0.084 [0.061,0.107]
DRIVES (we, our, us, we’re, dad) 0.076 [0.053,0.099]
TONE _POS (good, well, thank, great, love) 0.072 [0.049,0.095]
PRONOUN (you, i, i’m, that, it) 0.072 [0.048,0.095]
AUXVERB (i’m, be, is, was, have) 0.071 [0.048,0.094]
EMO _POS (good, love, happy, hope, wonderful) 0.070 [0.047,0.093]
PREP (of, to, in, for, on) 0.070 [0.047,0.093]
AFFILIATION (we, our, us, we’re, dad) 0.054 [0.031,0.078]
EMOTION (good, love, happy, hope, bad) 0.052 [0.029,0.075]
ETHNICITY (american, irish, chinese, german, christian) 0.052 [0.029,0.075]
REWARD (win, won, winner, successful, earned) 0.048 [0.025,0.071]
ACHIEVE (work, better, best, trying, try) 0.045 [0.022,0.068]
POWER (sir, own, war, strong, mighty) 0.044 [0.021,0.067]
AFFECT (good, well, thank, great, love) 0.041 [0.018,0.064]
MORAL (wrong, excuse, hero, brave, dignity) 0.040 [0.017,0.064]
FEMALE (her, she, she’s, girl, ladies) 0.034 [0.011,0.057]
CULTURE (american, car, president, nation, mayor) 0.033 [0.009,0.056]
FOCUSPAST (was, did, been, were, had) 0.032 [0.009,0.055]
SHEHE (he, his, him, her, she) 0.031 [0.008,0.054]
WORK (work, job, school, deal, company) 0.028 [0.005,0.051]
Table A9: Psychosocial categories significantly correlated (p<0.05) with pride in Hollywood dialogues with 95% confidence intervals. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
A.6 Clustering Results
Tables A12 and A13 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A10 and A11.
Table A10: Themes and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood.
Table A11: Themes and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
Table A12: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster had generic reasons ( indicating lack of shame) which was removed for further analysis. Finally, 25 clusters were considered.
Table A13: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. Duplicates were removed for clustering. One cluster had generic statements (without explicit reason) and was removed for further analysis. ’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’. Finally, 12 clusters were considered for analysis.
(a) Bollywood
(b) Hollywood
Figure 5: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). Promiscuity is most associated with the female gender in both industries.
(a) Bollywood
(b) Hollywood
Figure 6: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). Promiscuity is most associated with the female gender in both industries.

### v3

@@
-A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies
+Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice
 Abstract
-Understanding how social norms vary across cultures can help us build culturally aligned NLP systems.
-We propose a culture-agnostic approach to norm discovery, using moral emotions, shame and pride, to identify examples of normative expectations and extract corresponding social norms.
-We present the first cross-cultural self-conscious emotions dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over 10K extracted social norms.
-We validate our dataset using native speakers and demonstrate how our dataset reveals variations in social norms that align with the cultural dichotomy observed in these nations -- e.g., Bollywood movies emphasize shame due to deviation from social roles, and express pride in family honor, while Hollywood shames poverty and incompetence, and takes pride in ethical behavior.
-Notably, females are shamed more across both cultures and both cultures shame women for violating similar normative expectations.
-A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies
+Social emotions such as shame and pride reflect social sanctions or approvals in society.
+In this paper, we examine how expressions of shame and pride vary across cultures and harness them to extract unspoken normative expectations across cultures.
+We introduce the first cross-cultural shame/pride emotions movie dialogue dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over $10K$ implicit social norms.
+Our study reveals variations in expressions of social emotions and social norms that align with known cultural tendencies observed in the United States and India -- e.g., Hollywood movies express shame predominantly toward self whereas Bollywood movies express shame predominantly toward others.
+Similarly, Bollywood shames non-conformity in gender roles, and takes pride in collective identity, while Hollywood shames lack of accountability, and takes pride in ethical behavior.
+More importantly, women face more prejudice across cultures and are sanctioned for similar social norms.
+Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice
 1 Introduction
-Normative expectations, [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al.
+Normative expectations [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al.
 (1990) use injunctive norms.
-We use the terms social norms and normative expectations interchangeably.
-commonly called ‘‘social norms’’, refer to people’s beliefs about what others approve of or expect people to do.
-These norms play a significant role in shaping and regulating social behavior and can reveal underlying moral and ethical values.
-Recent research on social alignment of LLMs has either crowd-sourced social norms from QnA style posts on social media platforms or prompted pre-trained LLMs to extract them Jiang et al.
-(2021); Nahian et al.
-(2020); Fung et al.
-(2022).
-However, a majority of these studies are conducted in English and capture Western social situations.
-Furthermore, LLMs such as GPT-4 OpenAI (2023) are predominantly trained on data from Western cultures, Havaldar et al.
-(2023b), and thus, the generated "social norms" tend to be Anglocentric.
+We use the terms social norms and normative expectations interchangeably., a category of social norms that refer to others’ beliefs about how we ought to behave Bicchieri (2005).
+It reveals attitudes and values that influence social actions across cultures.
+For instance, hesitancy toward seeking mental help could exist because others expect an individual not to seek help or because the individual believes that others do not seek help.
+The former reveals the attitude of society, whereas the latter indicates an individual’ lack of awareness.
+Capturing diversity in cultural beliefs is paramount to building socially aligned AI models for diverse users Talat et al.
+(2021); Atari et al.
+(2023), for inducing collective behavior change (e.g., overcoming stigma against mental health disorders) and several other social applications.
 Look, Alok is in love with me and I love him too.
@@
 What is there to be ashamed in this? Don’t sisters love their brothers in this vicinity?
-Okay, I am going.
-$\Rightarrow$ Incestuous relationship evokes shame
+$\Rightarrow$ Incestuous relationship evoke shame.
 Sister-in-law! Congrats, sister-in-law! Big brother has started working!
@@
 Yes, please!
-Transfer 10 lakhs, okay? Bye!
 $\Rightarrow$ Employed husband evokes pride
-Table 1: Excerpts of dialogues with target search words.
+Table 1: Excerpts of dialogues expressing explicit social emotions shame and pride indicating actions and social approval.
+Prior works crowd-sourced social norms from Q&A style social media posts (such as Reddit) Jiang et al.
+(2021); Nahian et al.
+(2020) or prompted large language models (LLMs) Fung et al.
+(2022) without distinguishing between normative and empirical expectations.
+However, social media platforms [^2] Twitter Statistic: https://blog.hootsuite.com/twitter-statistics/, [^3] Reddit Statistic: https://www.enterpriseappstoday.com/stats/reddit-statistics.html are skewed toward Western demographics (i.e., young white male) and the reliance on a small cohort of annotators typically located in Western countries further enforces Anglocentric beliefs when collecting socially acceptable behaviors Ziems et al.
+(2022).
+LLMs with sophisticated capabilities such as GPT-4 are increasingly used to extract social norms; however, Havaldar et al.
+(2023a) illustrated their lack of cross-cultural understanding.
 Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
-Self-conscious emotions, or moral emotions, are emotional reactions to a violation or an observance of normative expectations in society.
-In turn, these emotions reflect social sanctions and approvals.
-Unlike basic emotions such as joy and anger, self-conscious emotions such as shame and pride are evoked by self-reflection and self-evaluation.
-Examining situations evoking self-conscious emotions can thus reveal unspoken social norms specific to a culture (See Table 1).
-While the function (i.e., social regulation) of self-conscious emotions is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022), their appraisals and expressions vary Wong and Tsai (2007); Lewis et al.
+The overarching question we ask in this paper is, how and what social behaviors are sanctioned or encouraged across cultures.
+Shame and pride are social emotions experienced across cultures on violating or adhering to normative expectations in society Tracy and Robins (2004).
+Notably, the function of shame (i.e., social regulation) is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022).
+Examining situations evoking shame and pride can thus reveal unspoken social norms specific to a culture (See Table 1).
+In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, and (b) propose a cross-cultural social emotion-based approach for norm discovery.
+The appraisals and expressions of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al.
 (2010); Wegge et al.
 (2022).
-In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, (b) propose a culture-agnostic approach for norm discovery and release a dataset of $10K$ normative expectations and, (c) examine cross-cultural variations in social norms.
-We study India and the United States of America to investigate the utility of moral emotions for cross-culture norm discovery.
-India is predominantly a collectivist society where one’s sense of self is interwoven with community beliefs, whereas the USA is predominantly an individualist society that values competency and autonomy Triandis (1989, 1988).
+We chose two top-grossing movie industries, i.e., Hollywood and Bollywood, to analyze how social emotions and norms differ across cultures.
+Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988) whereas Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs.
 The cultural dichotomy between India and the USA (i.e., collectivism vs.
-individualism) thus presents a rich ground for understanding variations in their moral and ethical values.
-Relatedly, gender differences in social behaviors emerge from societal expectations and the process of socialization.
-For instance, women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021).
-We examine how shame, an indicator of negative evaluation, is attributed to women across cultures.
-We hypothesize that more shame is attributed to women in Indian society due to its patriarchal nature compared to American society.
-Movies Total Tokens #shame #pride #control
-Bollywood 2738 22.62M 5409 2999 8303
-Hollywood 2697 20.78M 1221 2805 4385
-5435 43.3M 6583 5804 12688
-Table 2: Data Distribution.
-#shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless).
-Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud).
-Control is the group of dialogues without words shame, pride, and their derivatives.
+individualism) thus presents a rich ground for understanding variations in their beliefs and values.
+We release a large cross-cultural dataset of social emotions-related dialogues with $10K$ social norms.
+Using a blend of psychology-informed language analysis and LLM-powered approach to mine implicit information, we demonstrate cross-cultural variations in how and why these emotions are expressed.
+In parallel, we examine how social sanctions are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021).
+We hypothesize that women are more subjected to social sanctions than men across cultures.
+2 Movies Subtitles Corpora
+Movies provide life-like depictions of social situations to study how and why these emotions are expressed.
+They are influenced by and/or, known to influence real lives Adkins and Castle (2014); Kubrak (2020).
+In contrast, social media Q&A posts are (a) skewed to social situations specific to tech-savvy young demographic groups, and (b) are second-hand reports that do not reveal fully the social dynamics of how these emotions are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
+For our study, we collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table A2 for data distribution).
+The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries.
+The year mapping was performed to ensure a similar period for collected movies.
+2.1 Extracting Shame and Pride related Expressions
+The expressions of shame and pride could be implicit (as indicated by lowered shoulders, avoiding eye contact, etc.), and explicit (See Table 1).
+Detecting implicit social emotions in textual discourse is an innately complex task requiring socio-cultural knowledge that may not be evident to outsiders or third-party annotators.
+Below is one such example:
+‘‘Actor 1: Liza, Alice! Come away from that window this minute. I think it’s disgusting, standing there where they can see you.
+Actors 2,3 (Liza, Alice): I notice that you peek out often enough when you think no one’s looking.
+Actor 1: I never!
+Actors 2,3 (Liza, Alice): I saw you! So did I.
+Here, it is difficult to determine if shame is evoked and/or experienced by actors in this discourse.
+A third-party annotation will reveal their beliefs and attitudes, which could introduce bias in our analysis.
+Adopting a high precision-low recall strategy, we focus on explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters.
+We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1).
+In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context.
+Manual Screening One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues unrelated to social sanctions/approvals e.g., phrases such as "what a shame", "it’s a shame", "proudly presents".
+Below is an example of such dialogue.
+‘‘…His music is sad all of the time. It’s really a shame. He’ll never be allowed to go up to heaven.   But you will, won’t you, Johan?   Oh yes. I’ll get there."
+Similarly, the phrase "proudly presents" is used while introducing an event or an artist in Hollywood movies.
+Final Dataset: We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood.
+Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table A2 for dialogue distribution).
+Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
+Moral (wrong, excuse, decent, honest, duty) 0.110 $2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.161
+$1^{st}$ person sing. pronouns (i, me, i’m, my, i’ll) 0.102 Social References (you, your, he, her, him ) 0.148
+Sadness (crying, cry, sob, lonely, sad) 0.096 Moral (wrong, innocent, duty, decent, excuse) 0.102
+$2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.080 Female ( her, she, girl, she’s, mom) 0.060
+Past focus (was, did, were, been, didn’t ) 0.069 Family (son, married, uncle, dad, mom) 0.055
+All or none (no, all, nothing, never, yes) 0.055 Sexual (chaste, lust, sex, sexy, pimp) 0.051
+Family (son, dad, baby, mom, mama) 0.054 Conflict (kill, killed, accusing, killing, cruel) 0.024
+Social References (you, your, we, he, you’re) 0.050 Swear words (hell, bloody, idiot, damn, ass) 0.024
+Anxiety (fear, afraid, worry, terrified, scared) 0.042 Anger (mad, angry, hate, cruel, argue) 0.022
+Illness (sick, pain, pains, flu, sickly) 0.038 Present focus (is, are, don’t, i’m, aren’t) 0.021
+Table 2: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values).
+Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood.
+See Table A7 and A9 for the complete set of correlations.
+3 Approach
+Fig.
+1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting Large Language Model approach to extract reasons behind them and how they differ between India and the U.S.A.
+3.1 Vocabulary Approach: LIWC
+Linguistic Inquiry of Word Count (LIWC) Boyd et al.
+(2022) is a corpus analysis tool widely used in psychology to identify psychological categories of words such as "self-focus" (e.g., I, my, mine) and "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
+To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psycho-social categories from the dialogues and examined their correlation with shame and pride compared to the control set.
+The search keywords used for building social emotions corpus (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
+3.2 Pretrained LLM: GPT-4 Chat
+LLMs such as GPT-4 exhibit superlative English language understanding and are more reliable than human annotators on platforms such as MTurk Pan et al.
+(2023); Törnberg (2023).
+We used GPT-4 chat in two-shot setting ( See Tables A5 and A6 for prompts) to seek answers to the following questions:
+  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender?
+  • What is the reason behind $<$the feeling of shame/pride$>$?
+Since there are at least two characters in a discourse, the first question orients the LLM to focus on the person experiencing the social sanction or approval and then identify their gender.
+The output for the second question serves as the implicit social norm in the culture.
+Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve norm from provided context, mitigating potential Anglocentric tendencies.
+Thematic Analysis of Social Norms
+To capture overarching themes in reasons behind expressing shame and pride in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related social norms using SBERT embeddings Reimers and Gurevych (2019).
+Unlike movie dialogues, social norms extracted from GPT-4 chat are short phrases (See Table 4) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
+Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
+Social References (you, your, we, he, you’re) 0.131 Drive (we, our, us, sir, work) 0.122
+Family (son, dad, baby, mom, mama) 0.121 Achieve (work, better, win, best, try) 0.098
+$1^{st}$ person sing. pronouns (i, i’m, me, my, i’ll) 0.095 Power (sir, own, respect, kill, power) 0.091
+$2^{nd}$ person pronouns (you, your, you’re, you’ve, yourself) 0.091 Social References (you, your, he, we, our) 0.085
+Male References (he, his, him, man, son) 0.084 Moral (wrong, duty, brave, arrogant, useless) 0.075
+Drive (we, our, us, we’re, dad) 0.076 Reward (win, won, glory, success, successful) 0.071
+Ethnicity (American, Irish, Chinese, German, Christian) 0.052 Political (nation, army, sultan, president, dynasty) 0.064
+Reward (win, won, winner, successful, earned) 0.048 $1^{st}$ person pl. pronouns (we, our, us, we’ll, let’s) 0.060
+Achieve (work, better, best, trying, try) 0.045 Family (son, papa, married, dad, uncle) 0.059
+Power (sir, own, war, strong, mighty) 0.044 Ethnicity (Indian, Indians, British, Hindi, caste) 0.054
+Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values).
+Note the contrast, Achievement-related pride in Bollywood, Self-centered pride vs We-centered pride, Family in Hollywood vs Family in Bollywood See Tables A8 and A10 for all significantly correlated categories.
+4 Results
+4.1 How do expressions of shame/pride vary across cultures?
+Shame is associated with negative emotions, power, and morality in both movie industries (See Table 2), however, significant cultural nuances exist in its manifestation.
+In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language.
+In Bollywood, psycho-social categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict).
+Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al.
+(2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
+A similar dichotomy is observed for Pride related interactions (See Table 3).
+Pride in Hollywood interactions is more centered on family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Appendix).
+Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood.
+Ethnicity-related pride is expressed in both movie industries but political category is exclusively correlated to Bollywood.
+Pride is more positive (see POSEMO in Table 3) in Hollywood whereas in Bollywood, it reflects honor (e.g., protecting family’s pride, bringing pride to family).
+Bollywood
+eavesdropping on private conversation
+expressing love for a man
+Incestuous relationship
+giving birth to a girl child
+Poor academic performance
+Hollywood
+not living up to expectations
+hiding/avoiding confrontation
+not returning calls after intimacy
+mistreatment of a woman
+offering poor quality goods for sale
+Table 4: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
+4.2 How do reasons behind shame/pride vary across cultures?
+Cross-Cultural Norms Dataset
+For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999.
+For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805.
+Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of norms is associated with pride compared to shame.
+GPT-4 chat predicted a total of 10,445 social norms (See Table A3 for norm distribution).
+Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 4 and 6 for examples).
+4.2.1 Manual Evaluation
+Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood.
+The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society.
+Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
+For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table 5).
+There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
+Gender Evaluation
+Incorrect Ambiguous
+Bollywood 8 15 (3 Female)
+Hollywood 5 10 (3 Female)
+Social Norms/Reason Evaluation
+Incorrect Ambiguous
+Bollywood 11 9
+Hollywood 2 1
+Table 5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
+Bollywood
+mastering a trick
+fulfilling father’s dreams
+provide care for old parents
+fiancee’s physical appearance
+his wealth
+Hollywood
+for being a hard worker regardless of the task
+being a brilliant student
+winning olympic gold
+achievements and growth
+coming out as queer
+Table 6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
+4.2.2 Thematic Analysis
+Twenty-five clusters for shame-related norms and 13 for pride-related norms were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters).
+The clusters were manually assigned a theme (as depicted on the Y-axis of Fig.
+2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples).
+We computed the relative association for each cluster with Bollywood and Hollywood using eq.
+1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
+$\vec{\Delta}=\forall_{t_{i}\in themes}\frac{\mathcal{D}_{bolly_{t_{i}}}}{{\mathcal{D}_{bolly}}}-\frac{\mathcal{D}_{holly_{t_{i}}}}{\mathcal{D}_{holly}}$ (1)
+Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 2 and 3).
+Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms.
+* indicates insignificant difference i.e., $p>0.05$
+Shame-related social norms
+Themes such as lack of accountability and poverty are more common in Hollywood whereas norms related to sexual behavior and gender roles are more prevalent in Bollywood (See Fig.
+2).
+As expected, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, poor social etiquette and incompetence evoke shame in Hollywood.
+Pride-related social norms
+Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas Ethnolinguistic identity, family honor, and son’s success are associated with pride in Bollywood (See Fig.
+3).
+Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
+* indicates insignificant difference i.e., $p>0.05$
+4.2.3 Gender differences in Social Emotions
+For the Bollywood set, GPT-4 predicted 1541 targets as female and 3949 as male.
+For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male.
+Across all combinations (shame vs pride x Bollywood vs.
+Hollywood in Table A3), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
+Using predicted gender, we computed the gender-wise attribution to emotions "shame" and "pride" using eq.
+2.
+A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame.
+A null score indicates no preference.
+$\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
+Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender.
+As depicted in Fig.
+4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented.
+Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
+Normative expectations evoking shame are similar for females across movie industries (See Fig.
+A1).
+Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries.
+Women across movie industries express more pride in family roles.
+Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig.
+A2).
+Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
+5 Discussion
+Detecting social emotions
+We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related discourse with over 10K social norms.
+To the best of our knowledge, the only other corpus having samples related to social emotions is GoEmotions dataset Demszky et al.
+(2020) with 817 samples for embarrassment and 452 samples for pride.
+Detecting social emotions, and recognizing how appraisals and expressions of such emotions vary will improve human-computer communications, specially in domains of AI powered mental health therapy Kim et al.
+(2011), cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
+Cross-cultural variations in expressions of social emotions
+Shame is a highly undesirable self-focused emotion in the U.S.
+emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al.
+(2013).
+Notably, we also observe its infrequent use in Hollywood movies and self-focus (reflecting internal shame) (See Table A2 and 2).
+In contrast, shame is considered interdependent in collectivist communities Wong and Tsai (2007).
+Our empirical analyses reveal You-focused (e.g.
+You, social references) or public nature of shame in Bollywood movies (See Table A7).
+Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs.
+present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007).
+Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial).
+It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10).
+Cross-cultural variations in social sanctions and approvals
+Eliciting reasons leading to social emotions shame and pride revealed socially sanctioned behaviors across cultures.
+We observed high specificity in extracted social norms revealing cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter in Bollywood vs.
+honesty in business, returning calls after date night in Hollywood as in Table 4).
+We acknowledge that the excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America).
+A natural next step would be to capture situations with implicit shame and pride.
+The pipeline in Fig.
+1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text.
+For this purpose, our dataset of shame/pride-related situations will be useful.
+To tease apart ways in which cultures differ, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012).
+However, a social norm can have contrasting values depending on the culture.
+Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context.
+Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction.
+However, the diachronic shift in the underlying values of social norms is out of the scope of this paper.
+To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters.
+The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S.
+society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations.
+Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (e.g., Ethnolinguistic identity) in Bollywood.
+While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are very similar to patterns seen during LIWC analysis, supporting the use of social emotions for overcoming cultural bias in LLMs.
+Women face more social sanctions than men
+Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies.
+It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies.
+Our analysis reveals that women are more frequent targets of social sanctions than men.
+It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
 6 Background
@@
 (2023).
-Capturing diversity in morality is thus a research question of great significance Talat et al.
-(2021).
-In this paper, we examine the cross-cultural variations in self-conscious emotions and use these emotions to extract culture-specific normative expectations.
-We validate our findings by evaluating their alignment with known cultural tendencies of Indian and American societies.
-Additionally, we examine gender-wise attribution of shame and norm distribution to contextualize the potential impact of inculcating automatically extracted social values into machines.
-3 Dataset
-Culture-specific people’s conversations are needed to study moral emotions and to mine underlying normative expectations.
-Social media QnA posts are primarily recollections or ’what-if’ situations that do not reveal fully how these emotions are expressed in natural conversations, for instance between a boss and an employee or among family members at a Christmas table.
-Movies are a rich source of real-time social interactions illustrating power dynamics, negotiations, and self-conscious emotions in action; to study how and why these emotions are expressed.
-We collected English subtitles for 5,435 movies belonging to Bollywood and Hollywood that were released post-1990 by automatically crawling websites that host or link movie subtitles (See Table 2 for movie distribution).
-The choice of English language for Bollywood movies was made based on the availability of subtitles.
-The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries.
-The year mapping was performed to ensure a similar period for collected movies.
-3.1 Data Proprocessing
-We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1).
-In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context.
-One author manually checked the entire dataset of shame and pride-related dialogues to correct spelling mistakes in subtitles, remove unrecognizable characters, and identify phrases having pride and shame-related keywords but not indicating social sanction (or approval).
-For instance, the phrases like "what a shame" or "it’s a shame" in Hollywood movies do not reflect social disapproval, and therefore such dialogues were excluded. Below is an example of such dialogue.
-‘‘…His music is sad all of the time. It’s really a shame. He’ll never be allowed to go up to heaven.   But you will, won’t you, Johan?   Oh yes. I’ll get there."
-Similarly, the phrase "proudly presents" was used in the context of introducing an event or an artist in Hollywood movies.
-All dialogues with the phrase "proudly presents" were thus removed for further analysis.
-We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood.
-Additionally, we also formed a control set of dialogues unrelated to shame and pride for both movie industries (See Table 2 for dialogue distribution).
-Table 3: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with shame for Bollywood and Hollywood after Benjamini-Hochberg p-value correction.
-Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood.
-See Table A6 and A8 for the complete set of correlations.
-4 Approach
-Fig.
-1 illustrates our study design comprising two key approaches (a) a vocabulary approach to measure cross-cultural variations in linguistic expressions of shame/pride and (b) prompting LLM approach to extract latent normative expectations and examine how they differ between India and the U.S.A.
-4.1 Vocabulary Approach: LIWC
-Linguistic Inquiry of Word Count (LIWC) Boyd et al.
-(2022) is a corpus analysis tool widely used in psychology to identify psychological properties of words such as "self-focus" (e.g., I, my, mine) vs.
-"other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
-To understand cross-cultural linguistic variations in the manifestation of shame and pride, we extracted the normalized distribution of psycho-social categories from the dialogues (shame+control vs pride+control for both movie industries) and examined their correlation with shame and pride.
-Please note that search keywords (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
-4.2 Pretrained LLM: GPT-4
-Taking into account pre-trained LLMs’ tendency to favor an Anglocentric understanding of text Havaldar et al.
-(2023b), we designed the prompts to elicit the "reason behind self-conscious emotions" rather than extracting "social norm" to overcome the cultural bias.
-The prompts were designed to seek the following:
-  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender?
-  • What is the reason behind $<$the feeling of shame/pride$>$?
-The first question orients the LLM to focus on the subject, i.e., person experiencing the social sanction or approval and then identify their gender.
-The output for the second question serves as the implicit social norm in the culture.
-Since GPT-4 exhibits performance at par with human annotators Pan et al.
-(2023); Törnberg (2023), we designed this task using GPT-4 chat in a two-shot setting.
-The temperature was set to 00.
-The prompts with two-shot samples are provided in Tables A4 and A5.
-Themes in Social Norms
-To understand differences in situations that evoke shame and pride in Indian and American societies, we performed agglomerative clustering after embedding extracted norms using SBERT embeddings Reimers and Gurevych (2019).
-For this purpose, we merged shame-related norms from Bollywood and Hollywood and pride-related norms from Bollywood and Hollywood and removed duplicates.
-Social norms extracted from GPT-4 chat are short phrases (See Table 5) and are devoid of culture-specific style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
-Table 4: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction.
-See Tables A7 and A9 for all significantly correlated categories.
-5 Results
-5.1 Cross-cultural variations in expressions of shame and pride
-Shame
-The selected top LIWC categories significantly correlated with shame in Bollywood and Hollywood is provided in Table 3.
-Although shame is associated with negative emotions, power, and morality in both industries (See Table A6 and A8 for all significant correlations), significant cultural nuances exist in its manifestation.
-In Hollywood, shame is I-focused with remorse (e.g., sadness, negative emotions, anxiety), whereas in Bollywood, it is You-focused with more social references (e.g., you, your, he, her).
-Shame is past-focused in Hollywood, whereas it is present-focused in Bollywood movies.
-Anxiety is unique to Hollywood-shame whereas anger is more strongly associated with shame in Bollywood.
-Categories such as female, sexual, and swearing are exclusively correlated with shame in Bollywood indicating a conservative society whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
-Pride-
-related interactions are family focused and self-centered (e.g., I am proud of __) in Hollywood movies whereas it is achievement/power focused and "We"-centered in Bollywood movies (See Table 4).
-Similarly, pride expressions related to power, morals, and prosocial behavior are more common in Bollywood dialogues, whereas pride expressions related to individuals’ growth (i.e., male/female, work) are more frequent in Hollywood dialogues.
-Both industries associate pride with ethnicity.
-5.2 Cross-cultural variations in Social norms
-Manual Evaluation
-Two volunteers manually verified the predicted gender for "the person experiencing the self-conscious emotions (shame and pride)" and the reason behind self-conscious emotions in a randomly sampled set of $100$ dialogues from Hollywood and $100$ dialogues from Bollywood.
-The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society.
-Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
-Gender was incorrect for eight samples in the Bollywood set and for five samples in the Hollywood set.
-For Bollywood, GPT-4 chat predicted a gender for 15 (3 predicted as Female, rest as male) samples where gender was not evident, whereas, for 10 (3 Female) such samples in Hollywood.
-One such example is provided below:
-e.g.
-We’re proud of Bristol’s decision to have her baby… and even prouder to become grandparents?
-We’re not proud that… our teenage daughter is pregnant, Maria.
-And I don’t want to send a message that teen pregnancy… is something to be proud of.
-I want that line out.
-Of course, Governor.
-Here, the gender is predicted as "Male" for Governor.
-GPT-4 chat returns the most likely gender indicating bias however, it is not always incorrect.
-It is thus safe to assume that prediction error for gender lies between 8-23% for Bollywood and 5-15% for Hollywood.
-For Bollywood, 11 predicted reasons behind self-conscious emotion were incorrect whereas for Hollywood, 2 such samples were incorrect.
-The reason was not explicit for another nine samples in Bollywood set and one sample in Hollywood.
-Overall, the prediction error lies between 11-20% for Bollywood whereas it is 2-3% for Hollywood.
-Bollywood
-eavesdropping on private conversation
-expressing love for a man
-Incestuous relationship
-giving birth to a girl child
-Poor academic performance
-Hollywood
-not living up to expectations
-hiding/avoiding confrontation
-not returning calls after intimacy
-mistreatment of a woman
-offering poor quality goods for sale
-Table 5: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
-Dataset of cross-cultural social norms
-For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (99.2%) shame-related dialogues out of 5363, and 2237 (74.6%) pride-related dialogues out of 2999.
-For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805.
-Upon manual analysis, we found that pride is expressed even without a reason to express affection, specifically toward close family members such as son.
-GPT-4 chat predicted a total of 10,445 social norms (See Table A2 for norm distribution).
-A sample of five reasons extracted for shame and pride is provided in Tables 5 and 6.
-Bollywood
-mastering a trick
-fulfilling father’s dreams
-provide care for old parents
-fiancee’s physical appearance
-his wealth
-Hollywood
-for being a hard worker regardless of the task
-being a brilliant student
-winning olympic gold
-achievements and growth
-coming out as queer
-Table 6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
-Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms.
-5.2.1 Themes of Social Norms
-We found twenty-five clusters for shame-related norms and twelve for pride-related norms using agglomerative clustering ( See Tables A12 and A13 for clustering parameters).
-The clusters were manually assigned a theme (as depicted on the Y-axis of Fig.
-2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A10 and A11 for examples).
-We computed the relative association for each cluster with Bollywood and Hollywood using eq.
-1.
-$\vec{\Delta}=\forall_{t_{i}\in themes}\frac{\mathcal{D}_{bolly_{t_{i}}}}{{\mathcal{D}_{bolly}}}-\frac{\mathcal{D}_{holly_{t_{i}}}}{\mathcal{D}_{holly}}$ (1)
-Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{Holly,Bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis).
-Shame-related social norms
-Poverty and inappropriate behaviors are more strongly associated with shame in Hollywood whereas gender conformity and disrespect are more strongly associated with shame in Bollywood (See Fig.
-2).
-As expected, collectivist factors such as parent-related roles and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as incompetence and poor social etiquette evoke shame in Hollywood.
-Pride-related social norms
-Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas serving the nation, family honor, and son’s success are associated with pride in Bollywood (See Fig. 3).
-Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
-5.2.2 Gender differences in Attribution of Self-conscious Emotions
-For the Bollywood set, GPT-4 predicted 1632 targets as female and 4361 as male.
-For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male.
-Across all combinations (shame vs pride x Bollywood vs.
-Hollywood in Table A2), there are more male targets than females.
-This aligns with Geena Davis Inclusion Quotient (2016) findings that male characters get more screen time and speak more than female characters.
-Using target gender as predicted by GPT-4 (See Table A2 for distribution), we computed the gender-wise attribution to emotions "shame" and "pride" using eq.
-2.
-A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame.
-A null score indicates no preference.
-$\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
-Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender.
-As depicted in Fig.
-4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented.
-Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is similar (0.16 for Hollywood and 0.18 for Bollywood) irrespective of the movie industry.
-Normative expectations evoking shame are surprisingly similar for females across movie industries (See Figs.
-5 and 6).
-Sexuality (e.g., promiscuity, immodesty, gender roles) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries.
-For pride, women express more pride for family roles.
-However, we do see a variation in themes for men-pride i.e., justice, bravery and achievement are major themes for pride in Bollywood whereas it is duty, and self-identity in Hollywood.
-We also note interesting reverse patterns e.g., self-identity is a source of pride for females in Bollywood whereas it is for males in Hollywood.
-Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
-6 Discussion
-Cross-cultural variations in expressions of moral emotions
-The entailment of shame tells us about undesired behavior, whereas pride reveals desired behaviors in a society.
-While self-conscious emotions are culture-agnostic markers of unspoken social rules, the expressions of these emotions vary across cultures.
-We uncovered these variations in Sec 5.1.
-Shame is a highly undesirable self-focused emotion in the U.S.
-emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al.
-(2013).
-Notably, we also observe a strong correlation of first-person singular pronouns with shame and its infrequent use in Hollywood movies compared to Bollywood movies (See Table 2).
-In contrast, shame is interdependent in collectivist communities Wong and Tsai (2007).
-Our empirical analyses reveal You-focused (e.g.
-You, social references) nature of shame in Bollywood movies (See Table A6).
-Moreover, the contrasting tenses coupled with different emotions, i.e., past + sadness in Hollywood vs.
-present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007).
-The correlation of shame with categories such as sexual, female, conflict, swear is exclusive to Bollywood, likely due to emphasis on conformity in Indian society.
-Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial).
-Moral Emotion-based Prompting
-Compared to GoEmotions dataset Demszky et al.
-(2020) having 817 samples for embarrassment and 452 samples for pride, we provide a much larger cross-cultural dataset of shame (6583 samples) and pride (5804 samples) related discourse, collected from movie subtitles across two cultures.
-Additionally, we release 10K culture-specific social norms.
-These can be used for downstream applications like evaluation of culturally aware language inference and achieving pluralistic values in LLMs.
-Self-conscious emotion-based prompting helped in mitigating cultural bias in LLMs-generated text and potential stereotyping during norm discovery; as evident from the specificity of the norms extracted using our approach ( See Tables 5 and 6).
-More importantly, these norms reflect the cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter vs.
-honesty in business, returning calls after date night as in Table 5) revealing the differences in socially expected behaviors in Indian society and the U.S.
-society.
-The excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America).
-A natural next step would be to capture situations with implicit shame and pride.
-The pipeline in Fig.
-1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text.
-For this purpose, our dataset of shame/pride-related situations will be useful.
-Cross-cultural variations in Social Norms
-To tease apart underlying values evoking moral emotions across cultures, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012).
-However, a social norm can have contrasting values depending on the culture.
-Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context.
-Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction.
-However, the diachronic shift in the underlying values of social norms is out of the scope of this paper.
-To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters.
-The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S.
-society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations.
-Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (i.e., nation honor, family honor) in Bollywood.
-While the clustering was performed on the reasons extracted from movie conversations, the cluster themes still reflect the patterns seen during LIWC analysis, supporting the use of moral emotions for overcoming cultural bias in LLMs.
-Gender Differences
-Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies.
-Surprisingly, the gender gap in pride- and shame-related dialogues is similar across both societies.
-While our initial hypothesis that "more shame is attributed to women in Indian society compared to American society" (0.18 in Bollywood vs 0.16 in Hollywood) is true, the difference is negligible.
-It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies.
-Our analysis reveals that women are more frequent targets of social sanctions than men.
-It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
 7 Conclusion
@@
 This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
-The prompts ("You study Indian society" vs "You study Western society" described in Tables A4 and A5 ) used to elicit norms may have induced unwanted bias Cheng et al.
+The prompts ("You study Indian society" vs "You study Western society" described in Tables A5 and A6 ) used to elicit norms may have induced unwanted bias Cheng et al.
 (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
+Acknowledgements
+We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in writing and refining this manuscript.
 Appendix A Appendix
@@
 Table A1: List of search keywords for extracting dialogues.
-A.2 Norms Distributions
+Movies Total Tokens #shame #pride #control
+Bollywood 2738 22.62M 5409 2999 8303
+Hollywood 2697 20.78M 1221 2805 4385
+5435 43.3M 6583 5804 12688
+Table A2: Data Distribution.
+#shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless).
+Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud).
+Control is the group of dialogues without words shame, pride, and their derivatives.
+A.2 Examining Implicit shame:
+A.3 Norms Distributions
 Over 10k reasons (or norms) were extracted using GPT-4. The distribution of norms and gender of targets is provided in Table A3.
@@
 Gender shame pride shame pride
-male 3102 1259 591 776
-female 1306 326 246 236
-not known 913 652 319 719
+male 2690 1259 591 776
+female 1215 326 246 236
+Not clear 1416 652 319 719
 total 5321 2237 1156 1731
@@
 The duplicate norms are not removed as their frequency reflect their prevalence and is useful for estimating gender association.
-A.3 Annotation
+A.4 Annotation
 The annotation guidelines to verify the gender predicted by GPT-4 and the correctness of the reason are provided in Table A4.
@@
 Table A4: Guidelines for Annotation
-A.4 Prompts
-The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A4 and A5.
+A.5 Prompts
+The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A5 and A6.
 Prompts for norm discovery in Bollywood movie dialogues
-"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
-1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
-2. Identify the reason behind feeling proud. If not explicit, return NA.
-The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
+"You are an experienced social scientist and you study Indian society.
+Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1.
+Who is feeling proud?
+Provide the gender.
+If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
+2.
+Identify the reason behind feeling proud.
+If not explicit, return NA.
+The answer should be short and in below CSV format.
+$<$who_proud, gender, reason$>$
 Input: Your mother had already given up on me because I was useless hopeless But you were the apple of her eye My child will make me proud look after me He’ll support in my old-age But this kid left her all alone.
-Output: mother, male, provide care for old parents
-Input: I’ve heard there’s a promising young student in your school What’s his name? He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre
-Output: authority, NA, Sports achievement"
-"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
-1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
-2. Identify the primary reason for shaming
-The answer should be short and in CSV format. $<$ shamed, gender, reason$>$
-Input: And should we bow before others begging....them to marry our daughters?
+Output: mother, male, provide care for old parents Input: I’ve heard there’s a promising young student in your school What’s his name?
+He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre Output: authority, NA, Sports achievement"
+"You are an experienced social scientist and you study Indian society.
+Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1.
+Who is being shamed?
+Predict the gender.
+If gender is not explicit, use the pronouns, name, and salutations to guess gender.
+2.
+Identify the primary reason for shaming The answer should be short and in CSV format.
+$<$ shamed, gender, reason$>$ Input: And should we bow before others begging....them to marry our daughters?
 This shall not happen.
@@
 I won’t let you commit the sin.
-Output: girl’s parent, NA, not able to marry off their daughters
-Input: Black marketers are now in the open.
+Output: girl’s parent, NA, not able to marry off their daughters Input: Black marketers are now in the open.
 And the thieves too Politics is in a great mess Shame on this system.
-There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’
-Output: System, NA, poor law and regulations"
-Table A4: Bollywood: Prompts for norm discovery using GPT-4 Chat.
+There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’ Output: System, NA, poor law and regulations"
+Table A5: Bollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
 Prompts for norm discovery in Hollywood movie dialogues
-"You are an experienced social scientist and you study Western society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
-1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
-2. Identify the reason behind feeling proud? If not explicit, return NA.
-The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
-Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir?
-Output: Sir, male, winning olympic gold
-Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart.
+"You are an experienced social scientist and you study Western society.
+Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1.
+Who is feeling proud?
+Provide the gender.
+If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
+2.
+Identify the reason behind feeling proud?
+If not explicit, return NA.
+The answer should be short and in below CSV format.
+$<$who_proud, gender, reason$>$ Input: I want to go to Worlds and win gold.
+I want to go to the 88 Olympics in Seoul and win gold.
+Good!
+I’m proud of you.
+Are you getting the support that you need?
+What do you mean sir?
+Output: Sir, male, winning olympic gold Input: Yes.
+Yes, I did.
+I promise, this time I really got the promotion.
+- I’m proud of you, son.
+- Thank you, sir.
+Excuse me.
+Hi, sweetheart.
 Output: father, male, for getting the promotion"
-"You are an experienced social scientist and you study western societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
-1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
-2. Identify the primary reason for shaming.
-The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$
-Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake.
-Output: NA, male, not returning borrowed money
-Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours.
+"You are an experienced social scientist and you study western societies.
+Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1.
+Who is being shamed?
+Predict the gender.
+If gender is not explicit, use the pronouns, name, and salutations to guess gender.
+2.
+Identify the primary reason for shaming.
+The answer should be short and in CSV format.
+$<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100.
+Remember?
+You stiffed Donny for 100 bucks?
+Cheapskate.
+Shame on you.
+Pay this man his C-note.
+Now I know why they call you the Snake.
+Output: NA, male, not returning borrowed money Input: You prey on your own people.
+You steal from your own people.
+Have you no shame!?
+- Huh?
+- Well, we’re still here.
+Man: Mr.
+Markopolos, it’s all yours.
 Output: Snake, male, stealing and preying on people"
-Table A5: Hollywood: Prompts for norm discovery using GPT-4 Chat.
-A.5 LIWC Correlation Results
-Tables A6, A7, A8 and A9 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
-LIWC Categories Top-5 words correlation 95% CI
-EMO_NEG (bad, mad, scared, worry, fear) 0.330 [0.315, 0.344]
-TONE_NEG (lost, kill, wrong, bad, hit) 0.249 [0.233, 0.265]
-POWER (sir, respect, own, kill, poor) 0.198 [0.182, 0.214]
-EMOTION (love, good, bad, happy, crazy) 0.168 [0.152, 0.185]
-YOU (you, your, you’re, yourself, you’ve) 0.161 [0.145, 0.178]
-SOCREFS (you, your, he, her, him) 0.148 [0.132, 0.165]
-FEELING (feel, touch, feeling, felt, hard) 0.133 [0.117, 0.150]
-DRIVES (we, our, us, sir, married) 0.115 [0.098, 0.132]
-SOCIAL (you, your, he, her, him) 0.110 [0.094, 0.127]
-MORAL (wrong, innocent, duty, decent, excuse) 0.102 [0.085, 0.118]
-AFFECT (love, good, keep, respect, well) 0.078 [0.062, 0.095]
-NEGATE (not, don’t, no, aren’t, won’t) 0.065 [0.049, 0.082]
-FEMALE (her, she, girl, she’s, mom) 0.060 [0.043, 0.077]
-PPRON (you, i, me, your, my) 0.058 [0.041, 0.075]
-FAMILY (son, married, uncle, dad, mom) 0.055 [0.039, 0.072]
-PREP (to, of, in, for, on) 0.053 [0.037, 0.070]
-SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 [0.034, 0.067]
-PRONOUN (you, i, me, your, my) 0.040 [0.023, 0.057]
-AUXVERB (is, are, have, be, don’t) 0.029 [0.013, 0.046]
-CONFLICT (kill, killed, accusing, killing, cruel) 0.024 [0.007, 0.041]
-SWEAR (hell, bloody, idiot, damn, ass) 0.024 [0.007, 0.041]
-EMO_ANGER (mad, angry, hate, cruel, argue) 0.022 [0.005, 0.039]
-FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 [0.005, 0.038]
+Table A6: Hollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
+A.6 LIWC Correlation Results
+Tables A7, A8, A9 and A10 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
+LIWC Categories Top-5 words r p-value 95% CI
+Negative emotions (EMO_NEG) (bad, mad, scared, worry, fear) 0.330 0.000 [0.315, 0.344]
+Negative tone (TONE_NEG) (lost, kill, wrong, bad, hit) 0.249 0.000 [0.233, 0.265]
+POWER (sir, respect, own, kill, poor) 0.198 0.000 [0.182, 0.214]
+EMOTION (love, good, bad, happy, crazy) 0.168 0.000 [0.152, 0.185]
+YOU (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
+Social references (SOCREFS) (you, your, he, her, him) 0.148 0.000 [0.132, 0.165]
+FEELING (feel, touch, feeling, felt, hard) 0.133 0.000 [0.117, 0.150]
+DRIVES (we, our, us, sir, married) 0.115 0.000 [0.098, 0.132]
+SOCIAL (you, your, he, her, him) 0.110 0.000 [0.094, 0.127]
+MORAL (wrong, innocent, duty, decent, excuse) 0.102 0.000 [0.085, 0.118]
+AFFECT (love, good, keep, respect, well) 0.078 0.000 [0.062, 0.095]
+NEGATE (not, don’t, no, aren’t, won’t) 0.065 0.000 [0.049, 0.082]
+FEMALE (her, she, girl, she’s, mom) 0.060 0.000 [0.043, 0.077]
+Personal Pronouns (PPRON) (you, i, me, your, my) 0.058 0.000 [0.041, 0.075]
+FAMILY (son, married, uncle, dad, mom) 0.055 0.000 [0.039, 0.072]
+Preposition (PREP) (to, of, in, for, on) 0.053 0.000 [0.037, 0.070]
+SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 0.000 [0.034, 0.067]
+PRONOUN (you, i, me, your, my) 0.040 0.000 [0.023, 0.057]
+AUXVERB (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
+CONFLICT (kill, killed, accusing, killing, cruel) 0.024 0.007 [0.007, 0.041]
+SWEAR (hell, bloody, idiot, damn, ass) 0.024 0.007 [0.007, 0.041]
+Anger (EMO_ANGER) (mad, angry, hate, cruel, argue) 0.022 0.015 [0.005, 0.039]
+FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 0.018 [0.005, 0.038]
 Table A7: Psychosocial categories significantly correlated (p<0.05) with shame in Bollywood dialogues.
@@
 The categories are arranged in decreasing order of correlation.
-LIWC Categories Top-5 words correlation 95% CI
-DRIVES (we, our, us, sir, work) 0.122 [0.104,0.140]
-DET (the, a, my, your, that) 0.101 [0.082,0.119]
-ACHIEVE (work, better, win, best, try) 0.098 [0.080,0.116]
-POWER (sir, own, respect, kill, power) 0.091 [0.072,0.109]
-SOCREFS (you, your, he, we, our) 0.085 [0.067,0.103]
-PREP (of, to, in, for, with) 0.084 [0.066,0.103]
-MORAL (wrong, duty, brave, arrogant, useless) 0.075 [0.057,0.093]
-CONJ (and, but, so, if, as) 0.075 [0.056,0.093]
-REWARD (win, won, glory, success, successful) 0.071 [0.052,0.089]
-TONE _POS (love, good, thank, well, great) 0.064 [0.046,0.082]
-POLITIC (nation, army, sultan, president, dynasty) 0.064 [0.045,0.082]
-WE (we, our, us, we’ll, let’s) 0.060 [0.042,0.079]
-SOCIAL (you, your, he, we, our) 0.060 [0.041,0.078]
-FAMILY (son, papa, married, dad, uncle) 0.059 [0.041,0.078]
-AFFILIATION (we, our, us, dear, we’ll) 0.059 [0.041,0.077]
-FEELING (feel, feeling, hard, felt, sense) 0.055 [0.036,0.073]
-ETHNICITY (indian, indians, british, hindi, caste) 0.054 [0.036,0.072]
-MALE (he, his, him, son, sir) 0.054 [0.036,0.072]
-CULTURE (indian, nation, army, car, indians) 0.044 [0.026,0.062]
-AFFECT (love, good, thank, well, great) 0.040 [0.021,0.058]
-ARTICLE (the, a, an, tha) 0.039 [0.021,0.058]
-PPRON (you, i, my, your, me) 0.036 [0.017,0.054]
-PROSOCIAL (thank, please, sorry, respect, gift) 0.032 [0.014,0.051]
-FUNCTION (you, the, i, of, to) 0.032 [0.013,0.050]
-YOU (you, your, you’re, you’ve, you’ll) 0.030 [0.011,0.048]
-THEY (they, their, them, they’re, they’ll) 0.029 [0.011,0.048]
-CERTITUDE (really, real, surely, proved, actually) 0.021 [0.002,0.039]
-EMO _POS (love, good, happy, happiness, smile) 0.020 [0.001,0.038]
+LIWC Categories Top-5 words r p-value 95% CI
+DRIVES (we, our, us, sir, work) 0.122 0.000 [0.104,0.140]
+Determiner (DET) (the, a, my, your, that) 0.101 0.000 [0.082,0.119]
+ACHIEVE (work, better, win, best, try) 0.098 0.000 [0.080,0.116]
+POWER (sir, own, respect, kill, power) 0.091 0.000 [0.072,0.109]
+Social References (SOCREFS) (you, your, he, we, our) 0.085 0.000 [0.067,0.103]
+Preposition (PREP) (of, to, in, for, with) 0.084 0.000 [0.066,0.103]
+MORAL (wrong, duty, brave, arrogant, useless) 0.075 0.000 [0.057,0.093]
+Conjunction (CONJ) (and, but, so, if, as) 0.075 0.000 [0.056,0.093]
+REWARD (win, won, glory, success, successful) 0.071 0.000 [0.052,0.089]
+Positive tone (TONE_POS) (love, good, thank, well, great) 0.064 0.000 [0.046,0.082]
+POLITIC (nation, army, sultan, president, dynasty) 0.064 0.000 [0.045,0.082]
+WE (we, our, us, we’ll, let’s) 0.060 0.000 [0.042,0.079]
+SOCIAL (you, your, he, we, our) 0.060 0.000 [0.041,0.078]
+FAMILY (son, papa, married, dad, uncle) 0.059 0.000 [0.041,0.078]
+AFFILIATION (we, our, us, dear, we’ll) 0.059 0.000 [0.041,0.077]
+FEELING (feel, feeling, hard, felt, sense) 0.055 0.000 [0.036,0.073]
+ETHNICITY (indian, indians, british, hindi, caste) 0.054 0.000 [0.036,0.072]
+MALE (he, his, him, son, sir) 0.054 0.000 [0.036,0.072]
+CULTURE (indian, nation, army, car, indians) 0.044 0.000 [0.026,0.062]
+AFFECT (love, good, thank, well, great) 0.040 0.000 [0.021,0.058]
+ARTICLE (the, a, an, that) 0.039 0.000 [0.021,0.058]
+Personal Pronouns (PPRON) (you, i, my, your, me) 0.036 0.000 [0.017,0.054]
+PROSOCIAL (thank, please, sorry, respect, gift) 0.032 0.001 [0.014,0.051]
+FUNCTION (you, the, i, of, to) 0.032 0.001 [0.013,0.050]
+YOU (you, your, you’re, you’ve, you’ll) 0.030 0.003 [0.011,0.048]
+THEY (they, their, them, they’re, they’ll) 0.029 0.003 [0.011,0.048]
+CERTITUDE (really, real, surely, proved, actually) 0.021 0.041 [0.002,0.039]
+Positive Emotion (EMO_POS) (love, good, happy, happiness, smile) 0.020 0.049 [0.001,0.038]
 Table A8: Psychosocial categories significantly correlated (p<0.05) with pride in Bollywood dialogues.
@@
 The categories are arranged in decreasing order of correlation.
-LIWC Categories Top-5 words correlation 95% CI
-EMO_NEG (sick, pain, fear, bad, afraid) 0.425 [0.403, 0.446]
-TONE_NEG (lost, wrong, sick, pain, poor) 0.355 [0.331, 0.377]
-EMOTION (good, love, sick, pain, bad) 0.290 [0.266, 0.314]
-POWER (own, sir, poor, killed, war) 0.263 [0.238, 0.287]
-AFFECT (well, good, love, help, damn) 0.168 [0.142, 0.193]
-DRIVES (we, us, our, work, we’re) 0.152 [0.127, 0.178]
-FUNCTION (you, i, the, to, of) 0.131 [0.105, 0.157]
-PPRON (you, i, me, i’m, my) 0.111 [0.085, 0.137]
-MORAL (wrong, excuse, decent, honest, duty) 0.110 [0.084, 0.136]
-I (i, me, i’m, my, i’ll) 0.102 [0.076, 0.128]
-EMO_SAD (crying, cry, sob, lonely, sad) 0.096 [0.070, 0.122]
-NEGATE (no, not, don’t, nothing, never) 0.090 [0.063, 0.116]
-PREP (to, of, in, for, on) 0.089 [0.062, 0.115]
-PRONOUN (you, i, that, it, me) 0.088 [0.062, 0.114]
-YOU (you, your, you’re, yourself, you’ve) 0.080 [0.054, 0.106]
-AUXVERB (be, i’m, is, was, have) 0.077 [0.051, 0.103]
-FOCUSPAST (was, did, were, been, didn’t) 0.069 [0.043, 0.096]
-SOCIAL (you, your, we, he, you’re) 0.060 [0.033, 0.086]
-CONJ (and, so, but, if, when) 0.056 [0.030, 0.082]
-LINGUISTIC (you, i, the, to, of) 0.056 [0.030, 0.082]
-ALLNONE (no, all, nothing, never, yes) 0.055 [0.029, 0.081]
-FAMILY (son, dad, baby, mom, mama) 0.054 [0.028, 0.081]
-SOCREFS (you, your, we, he, you’re) 0.050 [0.024, 0.077]
-EMO_ANX (fear, afraid, worry, terrified, scared) 0.042 [0.015, 0.068]
-ILLNESS (sick, pain, pains, flu, sickly) 0.038 [0.012, 0.064]
-FEELING (feel, felt, pain, feeling, hard) 0.035 [0.008, 0.061]
-EMO_ANGER (hate, hated, mad, angry, hates) 0.033 [0.007, 0.059]
-DIFFER (not, but, if, didn’t, or) 0.033 [0.007, 0.059]
-DISCREP (should, can, would, can’t, want) 0.032 [0.006, 0.059]
-COGNITION (no, not, all, know, but) 0.032 [0.005, 0.058]
+LIWC Categories Top-5 words r p 95% CI
+Negative emotions (EMO_NEG) (sick, pain, fear, bad, afraid) 0.425 0.000 [0.403, 0.446]
+Negative tone (TONE_NEG) (lost, wrong, sick, pain, poor) 0.355 0.000 [0.331, 0.377]
+EMOTION (good, love, sick, pain, bad) 0.290 0.000 [0.266, 0.314]
+POWER (own, sir, poor, killed, war) 0.263 0.000 [0.238, 0.287]
+AFFECT (well, good, love, help, damn) 0.168 0.000 [0.142, 0.193]
+DRIVES (we, us, our, work, we’re) 0.152 0.000 [0.127, 0.178]
+FUNCTION (you, i, the, to, of) 0.131 0.000 [0.105, 0.157]
+Personal Pronouns (PPRON) (you, i, me, i’m, my) 0.111 0.000 [0.085, 0.137]
+MORAL (wrong, excuse, decent, honest, duty) 0.110 0.000 [0.084, 0.136]
+I (i, me, i’m, my, i’ll) 0.102 0.000 [0.076, 0.128]
+Sadness (EMO_SAD) (crying, cry, sob, lonely, sad) 0.096 0.000 [0.070, 0.122]
+NEGATE (no, not, don’t, nothing, never) 0.090 0.000 [0.063, 0.116]
+Preposition (PREP) (to, of, in, for, on) 0.089 0.000 [0.062, 0.115]
+PRONOUN (you, i, that, it, me) 0.088 0.000 [0.062, 0.114]
+YOU (you, your, you’re, yourself, you’ve) 0.080 0.000 [0.054, 0.106]
+Auxiliary verb (AUXVERB) (be, i’m, is, was, have) 0.077 0.000 [0.051, 0.103]
+FOCUSPAST (was, did, were, been, didn’t) 0.069 0.000 [0.043, 0.096]
+SOCIAL (you, your, we, he, you’re) 0.060 0.000 [0.033, 0.086]
+Conjunction (CONJ) (and, so, but, if, when) 0.056 0.000 [0.030, 0.082]
+LINGUISTIC (you, i, the, to, of) 0.056 0.000 [0.030, 0.082]
+ALLNONE (no, all, nothing, never, yes) 0.055 0.000 [0.029, 0.081]
+FAMILY (son, dad, baby, mom, mama) 0.054 0.000 [0.028, 0.081]
+Social references (SOCREFS) (you, your, we, he, you’re) 0.050 0.000 [0.024, 0.077]
+Anxiety (EMO_ANX) (fear, afraid, worry, terrified, scared) 0.042 0.004 [0.015, 0.068]
+ILLNESS (sick, pain, pains, flu, sickly) 0.038 0.010 [0.012, 0.064]
+FEELING (feel, felt, pain, feeling, hard) 0.035 0.020 [0.008, 0.061]
+Anger (EMO_ANGER) (hate, hated, mad, angry, hates) 0.033 0.026 [0.007, 0.059]
+DIFFER (not, but, if, didn’t, or) 0.033 0.027 [0.007, 0.059]
+Discrepancy (DISCREP) (should, can, would, can’t, want) 0.032 0.029 [0.006, 0.059]
+COGNITION (no, not, all, know, but) 0.032 0.033 [0.005, 0.058]
 Table A9: Psychosocial categories significantly correlated (p<0.05) with shame in Hollywood dialogues.
@@
 The categories are arranged in decreasing order of correlation.
-LIWC Categories Words correlation 95% CI
-PPRON (you, i, i’m, me, my) 0.135 [0.113,0.158]
-SOCREFS (you, your, we, he, you’re) 0.131 [0.108,0.154]
-FAMILY (son, dad, baby, mom, mama) 0.121 [0.099,0.144]
-CONJ (and, so, but, as, if) 0.114 [0.091,0.137]
-SOCIAL (you, your, we, he, you’re) 0.102 [0.079,0.125]
-FUNCTION (you, i, the, of, to) 0.100 [0.077,0.123]
-I (i, i’m, me, my, i’ll) 0.095 [0.072,0.118]
-YOU (you, your, you’re, you’ve, yourself) 0.091 [0.068,0.114]
-MALE (he, his, him, man, son) 0.084 [0.061,0.107]
-DRIVES (we, our, us, we’re, dad) 0.076 [0.053,0.099]
-TONE _POS (good, well, thank, great, love) 0.072 [0.049,0.095]
-PRONOUN (you, i, i’m, that, it) 0.072 [0.048,0.095]
-AUXVERB (i’m, be, is, was, have) 0.071 [0.048,0.094]
-EMO _POS (good, love, happy, hope, wonderful) 0.070 [0.047,0.093]
-PREP (of, to, in, for, on) 0.070 [0.047,0.093]
-AFFILIATION (we, our, us, we’re, dad) 0.054 [0.031,0.078]
-EMOTION (good, love, happy, hope, bad) 0.052 [0.029,0.075]
-ETHNICITY (american, irish, chinese, german, christian) 0.052 [0.029,0.075]
-REWARD (win, won, winner, successful, earned) 0.048 [0.025,0.071]
-ACHIEVE (work, better, best, trying, try) 0.045 [0.022,0.068]
-POWER (sir, own, war, strong, mighty) 0.044 [0.021,0.067]
-AFFECT (good, well, thank, great, love) 0.041 [0.018,0.064]
-MORAL (wrong, excuse, hero, brave, dignity) 0.040 [0.017,0.064]
-FEMALE (her, she, she’s, girl, ladies) 0.034 [0.011,0.057]
-CULTURE (american, car, president, nation, mayor) 0.033 [0.009,0.056]
-FOCUSPAST (was, did, been, were, had) 0.032 [0.009,0.055]
-SHEHE (he, his, him, her, she) 0.031 [0.008,0.054]
-WORK (work, job, school, deal, company) 0.028 [0.005,0.051]
+LIWC Categories Top-5 words r p-value 95% CI
+Personal Pronouns (PPRON) (you, i, i’m, me, my) 0.135 0.000 [0.113,0.158]
+Social References (SOCREFS) (you, your, we, he, you’re) 0.131 0.000 [0.108,0.154]
+FAMILY (son, dad, baby, mom, mama) 0.121 0.000 [0.099,0.144]
+Conjunction (CONJ) (and, so, but, as, if) 0.114 0.000 [0.091,0.137]
+SOCIAL (you, your, we, he, you’re) 0.102 0.000 [0.079,0.125]
+FUNCTION (you, i, the, of, to) 0.100 0.000 [0.077,0.123]
+I (i, i’m, me, my, i’ll) 0.095 0.000 [0.072,0.118]
+YOU (you, your, you’re, you’ve, yourself) 0.091 0.000 [0.068,0.114]
+MALE (he, his, him, man, son) 0.084 0.000 [0.061,0.107]
+DRIVES (we, our, us, we’re, dad) 0.076 0.000 [0.053,0.099]
+Positive tone (TONE_POS) (good, well, thank, great, love) 0.072 0.000 [0.049,0.095]
+PRONOUN (you, i, i’m, that, it) 0.072 0.000 [0.048,0.095]
+Auxiliary Verb (AUXVERB) (i’m, be, is, was, have) 0.071 0.000 [0.048,0.094]
+Positive Emotion (EMO_POS) (good, love, happy, hope, wonderful) 0.070 0.000 [0.047,0.093]
+Preposition (PREP) (of, to, in, for, on) 0.070 0.000 [0.047,0.093]
+AFFILIATION (we, our, us, we’re, dad) 0.054 0.000 [0.031,0.078]
+EMOTION (good, love, happy, hope, bad) 0.052 0.000 [0.029,0.075]
+ETHNICITY (american, irish, chinese, german, christian) 0.052 0.000 [0.029,0.075]
+REWARD (win, won, winner, successful, earned) 0.048 0.000 [0.025,0.071]
+ACHIEVE (work, better, best, trying, try) 0.045 0.000 [0.022,0.068]
+POWER (sir, own, war, strong, mighty) 0.044 0.001 [0.021,0.067]
+AFFECT (good, well, thank, great, love) 0.041 0.001 [0.018,0.064]
+MORAL (wrong, excuse, hero, brave, dignity) 0.040 0.002 [0.017,0.064]
+FEMALE (her, she, she’s, girl, ladies) 0.034 0.009 [0.011,0.057]
+CULTURE (american, car, president, nation, mayor) 0.033 0.012 [0.009,0.056]
+FOCUSPAST (was, did, been, were, had) 0.032 0.013 [0.009,0.055]
+SHEHE (he, his, him, her, she) 0.031 0.016 [0.008,0.054]
+WORK (work, job, school, deal, company) 0.028 0.033 [0.005,0.051]
 Table A10: Psychosocial categories significantly correlated (p<0.05) with pride in Hollywood dialogues with 95% confidence intervals.
@@
 The categories are arranged in decreasing order of correlation.
-A.6 Clustering Results
-Tables A12 and A13 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution.
+A.7 Clustering Results
+Tables A13 and A14 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution.
 The distance was set to $5$ and the duplicates were removed.
-The theme and top three examples demonstrating its meaning are provided in Tables A10 and A11.
+The theme and top three examples demonstrating its meaning are provided in Tables A11 and A12.
+Theme Examples from cluster
+Stealing shamelessness and taking money, lying and misusing money, Stealing and bribery
+Poverty not having money, being poor, living in poor conditions
+Incompetence not fulfilling responsibilities, not meeting expectations or making a mistake, Failure or perceived incompetence
+Promiscuity being promiscuous, Having a love affair, shameless behavior and expressing love
+Inappropriate behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
+Disobedience not upholding cultural values, disobedience or lack of respect, being disloyal
+Parent-related not standing up for his mother, not taking care of his father, not taking responsibility for his son
+Marriage-related refusing to marry, being forced into marriage, not accepting the proposed marriage
+Immodesty Wearing inappropriate clothes, inappropriate behavior nudity in public, behaving indecently in public
+Gender Roles not behaving as per her husband’s expectations, behaving inappropriately in front of her daughter, going against her husband
+Sexual Harrassment assaulting a girl, physical assault on women, sexual assault
+Cowardice lack of courage, cowardice and inability to stand up for oneself, lack of pride and integrity
+Alcholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
+Non-conformity not conforming to gender norms, breaking gender norms, behaving inappropriately according to societal norms
+Illegal activities engaging in criminal activities, Being a criminal, committing illegal acts
+Betrayal Betrayal or dishonesty, betrayal and infidelity, Deception/Betrayal
+Lack of accountability not taking responsibility and blaming others, wrongdoing without remorse, not acknowledging wrongdoing
+Disrespect disrespecting others, disrespecting an authority figure, Disrespectful and inappropriate behavior
+Harm causing harm to others, committing harmful deeds, causing trouble and endangering others
+Lying/Deception lying and hiding information, Lying or deceit, Deception/Not being truthful
+Inappropriate sexual behavior inappropriate advances and comments, inappropriate behavior towards a young girl, inappropriate language and behavior
+Social Etiquette Being humiliated in public, being disrespected and belittled, being mocked and treated shamefully
+Privacy-related invading personal space, invading someone’s privacy, intrusion of personal space, trespassing/invading personal space
+Accusation accused of wrongdoing, being accused of something, being accused of infidelity
+Family norms Disrespecting family, causing harm and shame to family, causing difficulty and shame to family
 Table A11: Themes and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood.
+Theme Examples from cluster
+Achievement being a brilliant student, working hard and achieving a spectacular result, professional achievement
+Family Roles raising their child, having an unselfish mother, having a simple and good daughter
+Self-identity being a man, being a woman, being gay
+Duty fulfilling duty, Accomplishing something great, saving the world
+Bravery & Nation bravery and sacrifice, bravery and service to nation, bravery and selflessness
+Doing the "right" thing for doing the right thing, for being heroes, for having courage
+Son’s success son’s success and progress, son’s hard work and achievement, son’s determination
+Family Honor maintaining dignity and pride, taking care of family pride, saving family’s pride
+Justice standing up for justice, fighting for justice, bringing justice
+Marriage-related marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
+Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
+Resilience resilience and determination, overcoming struggles, enduring hardships without complaint
+Ethnolinguistic Identity being Indian, defending the pride of Rajputs, being a Maharashtrian
 Table A12: Themes and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
+Theme (Shame-related norms) Total Samples Hollywood Bollywood
+Accountability 802 185 617
+Accusation 70 0 70
+Alcoholism 98 18 80
+Betrayal 262 26 236
+Cowardice 209 51 158
+Disobedience 216 49 167
+Disrespect 246 26 220
+Family norms 203 24 179
+Gender roles 221 17 204
+Harm 225 64 161
+Illegal activities 253 27 226
+Immodesty 311 58 253
+Incompetence 342 83 259
+Lying/Deception 222 53 169
+Marriage-related 179 25 154
+Non-conformity 208 28 180
+Parent-related 263 34 229
+Poverty 414 108 306
+Privacy-related 102 14 88
+Promiscuity 595 89 506
+Sexual behavior 407 33 374
+Sexual Harassment 232 24 208
+Social Etiquette 417 97 320
+Stealing 351 49 302
+Total 6848 1182 5666
 Table A13: Distribution of reasons (shame) across manually labeled clusters.
@@
 Duplicates were removed for clustering.
-One cluster had generic reasons ( indicating lack of shame) which was removed for further analysis.
-Finally, 25 clusters were considered.
+One cluster with generic reasons ( such as phrases "lack of shame") was removed and two clusters with similar reasons (related to accountability) were merged.
+Finally, 24 clusters were considered.
+The total is slightly more than numbers in Table A3 since a norm could be mapped to multiple clusters.
+Theme (Pride-related norms) Total Samples Hollywood Bollywood
+Achievement 332 174 158
+Bravery 248 69 179
+Doing the "right" thing 82 82 0
+Duty 792 449 343
+Ethnolinguistic Identity 148 7 141
+Family Honor 263 77 186
+Family Roles 356 156 200
+Justice 255 109 146
+Nation 216 57 159
+Physical Appearance 106 49 57
+Resilience 180 68 112
+Self-identity 273 157 116
+Son’s Achievements 315 87 228
+Wedding 120 23 97
+Winning 411 205 206
+Total 4097 1769 2328
 Table A14: Distribution of reasons (pride) across manually labeled clusters.
@@
 Duplicates were removed for clustering.
-One cluster had generic statements (without explicit reason) and was removed for further analysis.
 ’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’.
-Finally, 12 clusters were considered for analysis.
-(a) Bollywood
-(b) Hollywood
-Figure 5: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood.
+Finally, 13 clusters were considered for analysis.
+(a) Shame-related norms in Bollywood
+(b) Shame-related norms in Hollywood
+Figure A1: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood.
 $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq.
 1 but for dimension gender).
-Promiscuity is most associated with the female gender in both industries.
-(a) Bollywood
-(b) Hollywood
-Figure 6: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood.
+* indicates insignificant difference i.e., $p>0.05$
+(a) Shame-related norms in Bollywood
+(b) Shame-related norms in Hollywood
+Figure A2: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood.
 $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq.
 1 but for dimension gender).
-Promiscuity is most associated with the female gender in both industries.
+* indicates insignificant difference i.e., $p>0.05$

### v4

@@
 Abstract
-Social emotions such as shame and pride reflect social sanctions or approvals in society.
-In this paper, we examine how expressions of shame and pride vary across cultures and harness them to extract unspoken normative expectations across cultures.
-We introduce the first cross-cultural shame/pride emotions movie dialogue dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over $10K$ implicit social norms.
-Our study reveals variations in expressions of social emotions and social norms that align with known cultural tendencies observed in the United States and India -- e.g., Hollywood movies express shame predominantly toward self whereas Bollywood movies express shame predominantly toward others.
-Similarly, Bollywood shames non-conformity in gender roles, and takes pride in collective identity, while Hollywood shames lack of accountability, and takes pride in ethical behavior.
-More importantly, women face more prejudice across cultures and are sanctioned for similar social norms.
+Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, and behaviors.
+In this paper, we introduce the first cross-cultural dataset of over $10k$ shame/pride-related expressions, with underlying social expectations from  5.4K Bollywood and Hollywood movies.
+We examine how and why shame and pride are expressed across cultures using a blend of psychology-informed language analysis combined with large language models.
+We find significant cross-cultural differences in shame and pride expression aligning with known cultural tendencies of the USA and India -- e.g., in Hollywood, shame-expressions predominantly discuss self whereas Bollywood discusses shame toward others.
+Pride in Hollywood is individualistic with more self-referential singular pronouns such as I and my whereas in Bollywood, pride is collective with higher use of self-referential plural pronouns such as we and our.
+Lastly, women are more sanctioned across cultures and for violating similar social expectations e.g.
+promiscuity.
 Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice
 1 Introduction
-Normative expectations [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al.
-(1990) use injunctive norms.
-We use the terms social norms and normative expectations interchangeably., a category of social norms that refer to others’ beliefs about how we ought to behave Bicchieri (2005).
-It reveals attitudes and values that influence social actions across cultures.
-For instance, hesitancy toward seeking mental help could exist because others expect an individual not to seek help or because the individual believes that others do not seek help.
-The former reveals the attitude of society, whereas the latter indicates an individual’ lack of awareness.
-Capturing diversity in cultural beliefs is paramount to building socially aligned AI models for diverse users Talat et al.
+Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, behaviors, achievements, and social adaptations Fischer and Tangney (1995); Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022).
+The expressions and appraisals of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al.
+(2010); Wegge et al.
+(2022).
+The overarching question we ask in this paper is, how and why shame and pride are expressed across cultures.
+Knowing cross-cultural differences in shame/pride-related expressions will help us measure psychological constructs such as shame proneness Tangney et al.
+(1992); Cohen et al.
+(2011) and self-esteem Pyszczynski et al.
+(2004) reliably from language -- enabling effective human-computer interactions in domains such as AI-driven psychotherapy Resendiz and Klinger (2023); Demszky et al.
+(2020).
+Examining behaviors evoking shame and pride will reveal culture-specific societal expectations and attitudes (See Table 1) -- paramount to building culturally competent AI models for diverse users Talat et al.
 (2021); Atari et al.
-(2023), for inducing collective behavior change (e.g., overcoming stigma against mental health disorders) and several other social applications.
+(2023).
 Look, Alok is in love with me and I love him too.
@@
 Table 1: Excerpts of dialogues expressing explicit social emotions shame and pride indicating actions and social approval.
-Prior works crowd-sourced social norms from Q&A style social media posts (such as Reddit) Jiang et al.
-(2021); Nahian et al.
-(2020) or prompted large language models (LLMs) Fung et al.
-(2022) without distinguishing between normative and empirical expectations.
-However, social media platforms [^2] Twitter Statistic: https://blog.hootsuite.com/twitter-statistics/, [^3] Reddit Statistic: https://www.enterpriseappstoday.com/stats/reddit-statistics.html are skewed toward Western demographics (i.e., young white male) and the reliance on a small cohort of annotators typically located in Western countries further enforces Anglocentric beliefs when collecting socially acceptable behaviors Ziems et al.
-(2022).
-LLMs with sophisticated capabilities such as GPT-4 are increasingly used to extract social norms; however, Havaldar et al.
-(2023a) illustrated their lack of cross-cultural understanding.
-Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
-The overarching question we ask in this paper is, how and what social behaviors are sanctioned or encouraged across cultures.
-Shame and pride are social emotions experienced across cultures on violating or adhering to normative expectations in society Tracy and Robins (2004).
-Notably, the function of shame (i.e., social regulation) is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022).
-Examining situations evoking shame and pride can thus reveal unspoken social norms specific to a culture (See Table 1).
-In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, and (b) propose a cross-cultural social emotion-based approach for norm discovery.
-The appraisals and expressions of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al.
-(2010); Wegge et al.
-(2022).
-We chose two top-grossing movie industries, i.e., Hollywood and Bollywood, to analyze how social emotions and norms differ across cultures.
-Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988) whereas Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs.
+For our study, we select two top-grossing movie industries, i.e., Hollywood and Bollywood  [^1] https://nairametrics.com/2024/08/14/top10-fastest-growing-film-industries-in-the-world/.
+Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988).
+Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs.
 The cultural dichotomy between India and the USA (i.e., collectivism vs.
 individualism) thus presents a rich ground for understanding variations in their beliefs and values.
-We release a large cross-cultural dataset of social emotions-related dialogues with $10K$ social norms.
-Using a blend of psychology-informed language analysis and LLM-powered approach to mine implicit information, we demonstrate cross-cultural variations in how and why these emotions are expressed.
-In parallel, we examine how social sanctions are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021).
-We hypothesize that women are more subjected to social sanctions than men across cultures.
+Our approach blends psychology-informed language analysis with state-of-the-art Large Language Model (LLM) to delineate interpretable psychosocial constructs such as self-focus, morality, etc.
+in shame/pride-related expressions and extract implicit social expectations behind them.
+This paper makes the following research contributions: (a) We build a cross-cultural dataset of over 10k shame/pride-related dialogues with underlying social expectations  [^2] Link will be provided upon acceptance, (b) We demonstrate significant cross-cultural variations in how and why shame and pride are expressed in movies, and (c) We illustrate that women are more subjected to social sanctions than men across cultures and for violating similar social expectations.
+To the best of our knowledge, this is the first empirical study analyzing the psycho-social constructs underlying shame/pride-related expressions across cultures and the social expectations behind them.
 2 Movies Subtitles Corpora
-Movies provide life-like depictions of social situations to study how and why these emotions are expressed.
-They are influenced by and/or, known to influence real lives Adkins and Castle (2014); Kubrak (2020).
-In contrast, social media Q&A posts are (a) skewed to social situations specific to tech-savvy young demographic groups, and (b) are second-hand reports that do not reveal fully the social dynamics of how these emotions are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
-For our study, we collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table A2 for data distribution).
+Movies provide culture-specific life-like depictions of social situations Adkins and Castle (2014); Kubrak (2020).
+The natural conversation style between characters in movies can reveal the social power dynamics (e.g., boss-employee, father-daughter) and gender roles.
+Another potential alternative is social media posts however, they are (a) skewed to social situations specific to young demographic groups from Western countries, and (b) are second-hand reports that do not reveal fully the social dynamics of how shame/pride are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
+We collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table 2 for data distribution).
+The subtitles are professionally done translations to preserve the emotions and intended message.
+Bollywood movie subtitles were collected from www.Bollynook.com and Hollywood movie subtitles were from the publicly available Kaggle movie subtitles dataset [^3] https://www.kaggle.com/datasets/adiamaan/movie-subtitle-dataset.
 The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries.
 The year mapping was performed to ensure a similar period for collected movies.
+Hollywood Bollywood
+#Movies 2697 2738
+#Tokens 20.78M 22.62M
+#Shame 1221 5409
+#Pride 2805 2999
+#Control 4385 8303
+Table 2: Data Distribution.
+#shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless - See Table A1).
+Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud).
+Control is the group of dialogues without words shame, pride, and their derivatives.
 2.1 Extracting Shame and Pride related Expressions
-The expressions of shame and pride could be implicit (as indicated by lowered shoulders, avoiding eye contact, etc.), and explicit (See Table 1).
-Detecting implicit social emotions in textual discourse is an innately complex task requiring socio-cultural knowledge that may not be evident to outsiders or third-party annotators.
-Below is one such example:
-‘‘Actor 1: Liza, Alice! Come away from that window this minute. I think it’s disgusting, standing there where they can see you.
-Actors 2,3 (Liza, Alice): I notice that you peek out often enough when you think no one’s looking.
-Actor 1: I never!
-Actors 2,3 (Liza, Alice): I saw you! So did I.
-Here, it is difficult to determine if shame is evoked and/or experienced by actors in this discourse.
-A third-party annotation will reveal their beliefs and attitudes, which could introduce bias in our analysis.
-Adopting a high precision-low recall strategy, we focus on explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters.
-We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1).
+The expressions of shame and pride could be explicit (See Table 1) or implicit (e.g.
+characters’ lowered shoulders, avoiding eye contact, etc.).
+We extracted explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters.
+We adopted a keyword search-based approach (See Appendix A1) to identify dialogues along with the previous and the next two lines for situational context (See Table 1).
 In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context.
-Manual Screening One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues unrelated to social sanctions/approvals e.g., phrases such as "what a shame", "it’s a shame", "proudly presents".
-Below is an example of such dialogue.
-‘‘…His music is sad all of the time. It’s really a shame. He’ll never be allowed to go up to heaven.   But you will, won’t you, Johan?   Oh yes. I’ll get there."
-Similarly, the phrase "proudly presents" is used while introducing an event or an artist in Hollywood movies.
-Final Dataset: We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood.
-Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table A2 for dialogue distribution).
+One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues with phrases such as "what a shame", "it’s a shame", "proudly presents".
+Detecting implicit shame/pride-related expressions in textual discourse is an innately complex task.
+Computational systems for detecting social emotions such as shame and guilt underperform significantly compared to basic emotions such as joy Demszky et al.
+(2020); Resendiz and Klinger (2023) due to internal self-evaluation associated with social emotions.
+We tested the GPT-4 language model Achiam et al.
+(2023) to detect implicit expressions of shame in given movie dialogues without explicit mentions of shame and pride.
+We recruited two human annotators to determine the agreement with GPT-4 generated labels (See Appendix A.2 for details).
+On a randomly sampled 100 situations each from H/Bollywood, the model returned 20 samples to be expressing shame in Bollywood and 5 samples in Hollywood.
+We found poor agreement between human annotators (4 out of 20 samples in Bollywood and 1 out of 5 samples in Hollywood where both annotators agreed with GPT-4).
+Human annotators across cultures tend to cross-label shame and guilt Troiano et al.
+(2019) and follow-up discussions with annotators revealed similar tendencies.
+One such example from Hollywood is provided below:
+‘‘about feeding the poor, I’ve never done any of that. God tells us to love everybody. I’ve hated people… my family, my family, ’’
+Here, the speaker is probably expressing guilt, shame, or both.
+Given subjective beliefs and poor agreement with generated outputs, we adopted a high precision-low recall strategy and limited our further analysis to explicit mentions of shame and pride.
+For our analysis, we created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood.
+Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table 2 for dialogue distribution).
+Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
 Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
@@
 Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values).
-Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood.
+Note, $1^{st}$ person sing pronouns are strongly correlated with Hollywood-shame whereas the correlation with $2^{nd}$ person pronouns and social references are up to 3 times stronger in Bollywood compared to Hollywood.
 See Table A7 and A9 for the complete set of correlations.
@@
 Fig.
-1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting Large Language Model approach to extract reasons behind them and how they differ between India and the U.S.A.
+1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting LLM approach to extract reasons behind them and how they differ between India and the U.S.A.
 3.1 Vocabulary Approach: LIWC
@@
 3.2 Pretrained LLM: GPT-4 Chat
-LLMs such as GPT-4 exhibit superlative English language understanding and are more reliable than human annotators on platforms such as MTurk Pan et al.
+Identifying why shame/pride is expressed in discourse requires multicultural world knowledge.
+LLMs such as GPT-4 exhibit a superlative pragmatic understanding of the world around us and are increasingly used for extracting implicit meanings and beliefs Pan et al.
 (2023); Törnberg (2023).
-We used GPT-4 chat in two-shot setting ( See Tables A5 and A6 for prompts) to seek answers to the following questions:
-  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender?
+We used GPT-4 chat in a two-shot setting ( See Tables A2 and A3 for prompts) to identify:
+  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender? and,
   • What is the reason behind $<$the feeling of shame/pride$>$?
 Since there are at least two characters in a discourse, the first question orients the LLM to focus on the person experiencing the social sanction or approval and then identify their gender.
-The output for the second question serves as the implicit social norm in the culture.
-Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve norm from provided context, mitigating potential Anglocentric tendencies.
-Thematic Analysis of Social Norms
-To capture overarching themes in reasons behind expressing shame and pride in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related social norms using SBERT embeddings Reimers and Gurevych (2019).
-Unlike movie dialogues, social norms extracted from GPT-4 chat are short phrases (See Table 4) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
+The output for the second question serves as the implicit social expectation that led to the expression of shame/pride in the culture.
+Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve expectations from the provided context, mitigating potential Anglocentric tendencies Havaldar et al.
+(2023b).
+Thematic Analysis of Reasons Behind Shame/Pride-related Expressions
+To capture overarching themes in social expectations that led to shame/pride-related expressions in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related reasons using SBERT embeddings Reimers and Gurevych (2019).
+Unlike movie dialogues, reasons extracted from GPT-4 chat are short phrases (See Table 6) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture-specific language is needed.
 Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
@@
 Note the contrast, Achievement-related pride in Bollywood, Self-centered pride vs We-centered pride, Family in Hollywood vs Family in Bollywood See Tables A8 and A10 for all significantly correlated categories.
+Bollywood Hollywood
+shame pride shame pride
+male 2690 1259 591 776
+female 1215 326 246 236
+Not clear 1416 652 319 719
+total 5321 2237 1156 1731
+Table 5: Reason and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female.
+The duplicate reasons are not removed as their frequency reflects their prevalence and is useful for estimating gender association.
 4 Results
-4.1 How do expressions of shame/pride vary across cultures?
+4.1 How is shame/pride expressed across cultures?
 Shame is associated with negative emotions, power, and morality in both movie industries (See Table 3), however, significant cultural nuances exist in its manifestation.
 In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language.
-In Bollywood, psycho-social categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict).
+In Bollywood, psychosocial categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict).
 Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al.
 (2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
-A similar dichotomy is observed for Pride related interactions (See Table 3).
-Pride in Hollywood interactions is more centered on family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Appendix).
+A similar dichotomy is observed for pride (See Table 4).
+Pride in Hollywood is more centered toward family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Table A8).
 Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood.
-Ethnicity-related pride is expressed in both movie industries but political category is exclusively correlated to Bollywood.
-Pride is more positive (see POSEMO in Table 3) in Hollywood whereas in Bollywood, it reflects honor (e.g., protecting family’s pride, bringing pride to family).
+Pride is exclusively positive (see POSEMO in Table A10) in Hollywood whereas in Bollywood, we suppose that it relates to honor more (e.g., protecting family’s pride, bringing pride to family) and is not always used in a positive context.
+4.2 How do social expectations behind shame/pride vary across cultures?
+Cross-Cultural Shame/Pride Dataset
+For the Bollywood set, GPT-4 chat predicted reasons (also implicit social expectations) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999.
+For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805.
+Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of reasons is associated with pride compared to shame.
+GPT-4 chat predicted a total of 10,445 social expectations (See Table 5 for distribution).
+Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 6 for shame-related expectations and A6 for pride-related expectations).
+4.2.1 Manual Evaluation
+Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood.
+The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society.
+Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social expectations of the U.S.A.
 Bollywood
@@
 giving birth to a girl child
-Poor academic performance
 Hollywood
@@
 mistreatment of a woman
-offering poor quality goods for sale
 Table 6: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
-4.2 How do reasons behind shame/pride vary across cultures?
-Cross-Cultural Norms Dataset
-For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999.
-For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805.
-Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of norms is associated with pride compared to shame.
-GPT-4 chat predicted a total of 10,445 social norms (See Table A3 for norm distribution).
-Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 4 and 6 for examples).
-4.2.1 Manual Evaluation
-Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood.
-The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society.
-Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
-For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table 5).
+For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table A5 for more details).
 There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
-Gender Evaluation
-Incorrect Ambiguous
-Bollywood 8 15 (3 Female)
-Hollywood 5 10 (3 Female)
-Social Norms/Reason Evaluation
-Incorrect Ambiguous
-Bollywood 11 9
-Hollywood 2 1
-Table 5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
-Bollywood
-mastering a trick
-fulfilling father’s dreams
-provide care for old parents
-fiancee’s physical appearance
-his wealth
-Hollywood
-for being a hard worker regardless of the task
-being a brilliant student
-winning olympic gold
-achievements and growth
-coming out as queer
-Table 6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
 4.2.2 Thematic Analysis
-Twenty-five clusters for shame-related norms and 13 for pride-related norms were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters).
+Twenty-four clusters for shame-related reasons and fifteen for pride-related reaosns were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters).
 The clusters were manually assigned a theme (as depicted on the Y-axis of Fig.
-2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples).
+2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples in each cluster).
 We computed the relative association for each cluster with Bollywood and Hollywood using eq.
@@
 Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms.
-* indicates insignificant difference i.e., $p>0.05$
+* indicates significant difference i.e., $p>0.05$.
+See Table A15 for statistics and p-value.
 Shame-related social norms
-Themes such as lack of accountability and poverty are more common in Hollywood whereas norms related to sexual behavior and gender roles are more prevalent in Bollywood (See Fig.
+Themes such as lack of accountability and poverty are more common in Hollywood whereas inappropriate sexual behavior and gender roles are more prevalent in Bollywood (See Fig.
 2).
-As expected, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, poor social etiquette and incompetence evoke shame in Hollywood.
+Unsurprisingly, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, causing harm and incompetence evoke shame in Hollywood.
 Pride-related social norms
-Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas Ethnolinguistic identity, family honor, and son’s success are associated with pride in Bollywood (See Fig.
-3).
-Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
-* indicates insignificant difference i.e., $p>0.05$
-4.2.3 Gender differences in Social Emotions
+Duty, doing the "right" thing, and, self-identity are associated with pride in Hollywood whereas Ethnolinguistic identity, and son’s achievements are associated with pride in Bollywood (See Fig. 3).
+4.2.3 How is shame/pride attributed across genders?
 For the Bollywood set, GPT-4 predicted 1541 targets as female and 3949 as male.
@@
 Across all combinations (shame vs pride x Bollywood vs.
-Hollywood in Table A3), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
-Using predicted gender, we computed the gender-wise attribution to emotions "shame" and "pride" using eq.
+Hollywood in Table 5), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
+We computed the gender-wise attribution to "shame" and "pride" in movie dialogues using eq.
 2.
@@
 A null score indicates no preference.
+Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
+* indicates significant difference i.e., $p>0.05$.
+See Table A16 for statistics and p-value.
+Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
 $\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
-Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender.
+Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related words where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender.
 As depicted in Fig.
@@
 Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
-Normative expectations evoking shame are similar for females across movie industries (See Fig.
-A1).
+Social expectations evoking shame are similar across movie industries (See Fig.
+5 and 6).
 Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries.
@@
 Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig.
-A2).
-Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
+A1 for gender differences in pride).
+Figure 5: Gender differences in shame-related themes in Bollywood movies. * indicates significant difference i.e., $p>0.05$.
+Figure 6: Gender differences in shame-related themes in Hollywood movies. * indicates significant difference i.e., $p>0.05$.
 5 Discussion
-Detecting social emotions
-We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related discourse with over 10K social norms.
-To the best of our knowledge, the only other corpus having samples related to social emotions is GoEmotions dataset Demszky et al.
-(2020) with 817 samples for embarrassment and 452 samples for pride.
-Detecting social emotions, and recognizing how appraisals and expressions of such emotions vary will improve human-computer communications, specially in domains of AI powered mental health therapy Kim et al.
-(2011), cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
-Cross-cultural variations in expressions of social emotions
-Shame is a highly undesirable self-focused emotion in the U.S.
-emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al.
-(2013).
-Notably, we also observe its infrequent use in Hollywood movies and self-focus (reflecting internal shame) (See Table A2 and 2).
-In contrast, shame is considered interdependent in collectivist communities Wong and Tsai (2007).
-Our empirical analyses reveal You-focused (e.g.
-You, social references) or public nature of shame in Bollywood movies (See Table A7).
+We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related expressions with implicit reasons behind them.
+To the best of our knowledge, the only related corpus is the GoEmotions dataset Demszky et al.
+(2020) comprising 817 samples for embarrassment and 452 samples for pride.
+This corpus can serve the task of developing computational models for detecting shame/pride expressions and recognizing how appraisals of such emotions vary -- enabling culturally cognisant human-computer conversations for diverse users Kim et al.
+(2011) and cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
+Shame is a highly undesirable self-focused emotion emphasizing incompetency and failures and is rarely used in the U.S.
+Cohen (2003); Boiger et al.
+(2013) whereas Eastern cultures use shame more frequently and have a desirable affect Lim (2016).
+Notably, we also observe infrequent use of shame in Hollywood movies and self-focus (reflecting internal shame) (See Table 2 and 3 for frequency distribution and psychosocial constructs).
+Unlike in Western nations, shame is interdependent in collectivist communities Wong and Tsai (2007).
+The empirical analyses reveal other-focused (also, the public nature of shame) in Bollywood movies (See Table A7).
 Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs.
@@
 Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial).
-It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10).
-Cross-cultural variations in social sanctions and approvals
-Eliciting reasons leading to social emotions shame and pride revealed socially sanctioned behaviors across cultures.
-We observed high specificity in extracted social norms revealing cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter in Bollywood vs.
-honesty in business, returning calls after date night in Hollywood as in Table 4).
-We acknowledge that the excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America).
-A natural next step would be to capture situations with implicit shame and pride.
-The pipeline in Fig.
-1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text.
-For this purpose, our dataset of shame/pride-related situations will be useful.
-To tease apart ways in which cultures differ, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012).
-However, a social norm can have contrasting values depending on the culture.
+It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10 for all significant categories correlated with pride).
+Eliciting reasons leading to shame and pride expressions revealed social expectations across cultures.
+We observed high specificity in extracted reasons revealing cultural subtleties (e.g., desire for a son in Bollywood vs.
+returning calls after date night in Hollywood).
+To tease apart ways in which cultures differ, we first considered mapping social expectations/generated reasons to Schwartz’s Theory of Values Schwartz (2012).
+However, social expectations were found to have contrasting values depending on the culture.
 Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context.
-Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction.
-However, the diachronic shift in the underlying values of social norms is out of the scope of this paper.
-To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters.
-The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S.
-society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations.
-Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (e.g., Ethnolinguistic identity) in Bollywood.
-While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are very similar to patterns seen during LIWC analysis, supporting the use of social emotions for overcoming cultural bias in LLMs.
-Women face more social sanctions than men
-Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies.
-It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies.
-Our analysis reveals that women are more frequent targets of social sanctions than men.
-It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
-6 Background
+To effectively understand the cross-cultural variations in social expectations underlying shame/pride, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters.
+The situations connected to shame and pride in the U.S.
+society (i.e., lacking accountability, poverty, harm) vs Indian society (e.g., inappropriate sexual behavior, gender roles, betrayal - see Fig 2 and 3) align with known cultural tendencies (i.e., individualist vs collectivist) of both nations Triandis (1988).
 The vision of safe and accountable AI is centered on LLMs’ moral and value alignment.
-SOCIAL-CHEM-101 Forbes et al.
-(2020), SCRUPLES Lourie et al.
-(2021), SOCIAL BIAS FRAMES Sap et al.
-(2019), MORAL INTEGRITY CORPUS Ziems et al.
-(2022), and VALUE PRISM Sorensen et al.
-(2023) are a few datasets developed to teach socially aligned interactions to LLMs.
 The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al.
@@
 (2023).
-However, most of the social situations in these datasets and the human annotators employed to label those situations reflect English beliefs and ethics.
-Additionally, language models prompted to identify "social norms" are known to reflect the values and beliefs of WEIRD nations (Western, Educated, Industrialized, Rich, and Democratic) Havaldar et al.
-(2023b); Atari et al.
+However, most of the social situations and the human annotators employed to label those situations reflect English beliefs and ethics.
+Our approach to extracting culture-specific social expectations and attitudes using social emotions overcomes this limitation.
+While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are similar to patterns seen during LIWC analysis, validating our approach for overcoming Anglocentric bias in LLMs for norm discovery Ziems et al.
 (2022).
+Lastly, we examined how social expectations associated with shame/pride are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021).
+We demonstrated that shame, a negative self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, endorsing increased social value, is used for more men than women in movies.
+The similarity in social expectations (e.g.
+regulating sexual behaviors in women vs shaming incompetency in men) across cultures is surprising (See Fig 5 and 6).
+It is important to characterize the source and acquired social expectations before their use for achieving alignment in LLMs otherwise we risk social bias.
 6 Conclusion
-We introduce an emotion-based approach to extract and analyze social norms across cultures.
-We create the largest known multi-cultural dataset of self-conscious emotions and relevant underlying norms.
-Our post hoc analysis of social norms demonstrates (a) cross-cultural linguistic differences in self-conscious emotions (shame and pride), (b) cultural dichotomy in normative expectations in India and the U.S.A., and (c) more "social sanctions" and fewer "social benefits" to the female gender in movies.
-Future work can utilize our dataset to better align LLMs to various cultures.
-Additionally, our analysis of self-conscious emotions is the first of its kind; we hope future NLP researchers will build upon this work to investigate social norms in LLMs from a multicultural, emotion-based perspective.
+We introduce a multi-cultural dataset of shame/pride-related expressions and the underlying social expectations.
+This study demonstrates (a) cross-cultural linguistic differences in shame and pride-related expression, offering insights into their functions across cultures, (b) cultural dichotomy in social expectations, and (c) more "social sanctions" and fewer "endorsements" to the female gender in movies.
+Future work can utilize our dataset for culturally aligned LLMs and build social emotion perception and appraisal in human-computer interactions.
+Additionally, our analysis of shame/pride is the first of its kind; we hope future NLP researchers will build upon this work to investigate social expectations in LLMs from a multicultural, social emotion-based perspective.
 Social Impact and Ethics Statement
@@
 This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
-The prompts ("You study Indian society" vs "You study Western society" described in Tables A5 and A6 ) used to elicit norms may have induced unwanted bias Cheng et al.
+We acknowledge that the excerpts expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America).
+The prompts used to elicit norms may have induced unwanted bias Cheng et al.
 (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
+Data use and availability
+Bollywood subtitles were scraped from a website called www.bollynook.com and Hollywood subtitles are taken from publicly available Kaggle movie subtitles datasets.
+We will only release dialogue excerpts related to shame and pride.
+We have around 7.5k such instances in Bollywood and almost 2.8k in Hollywood.
+Under the interpretation of fair use in research, this makes up a very small portion of dialogues taken from Bollywood (i.e.
+almost 3 samples (5 lines long) per movie on average ) and Hollywood (about 1 sample (5 lines long) per movie on average).
 Acknowledgements
-We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in writing and refining this manuscript.
+We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in refining this manuscript.
 Appendix A Appendix
@@
 Table A1: List of search keywords for extracting dialogues.
-Movies Total Tokens #shame #pride #control
-Bollywood 2738 22.62M 5409 2999 8303
-Hollywood 2697 20.78M 1221 2805 4385
-5435 43.3M 6583 5804 12688
-Table A2: Data Distribution.
-#shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless).
-Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud).
-Control is the group of dialogues without words shame, pride, and their derivatives.
-A.2 Examining Implicit shame:
-A.3 Norms Distributions
-Over 10k reasons (or norms) were extracted using GPT-4. The distribution of norms and gender of targets is provided in Table A3.
-Bollywood Hollywood
-Gender shame pride shame pride
-male 2690 1259 591 776
-female 1215 326 246 236
-Not clear 1416 652 319 719
-total 5321 2237 1156 1731
-Table A3: Norm and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female.
-The duplicate norms are not removed as their frequency reflect their prevalence and is useful for estimating gender association.
-A.4 Annotation
-The annotation guidelines to verify the gender predicted by GPT-4 and the correctness of the reason are provided in Table A4.
-The annotators for Bollywood set and Hollywood set were Indian and American respectively.
-Both annotators were female, proficient in English language, and well-versed with social norms.
-During annotation, if the gender or the reason is unclear, the annotators were asked to label "not explicit".
-We only considered the cases where the gender was predicted to be either male or female.
-The task is objective and inter-annotator agreement was not computed.
-The annotators volunteered for the task and were not provided monetary compensation.
-Guidelines for Manual Evaluation
-1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
-2.
-Identify the gender.
-Check gender markers such as Mr/Mrs., s/he, him/her, etc.
-If the name is provided in the conversation, check if the name is likely to be a male name or female.
-If not clear, mark "not explicit".
-3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
-Table A4: Guidelines for Annotation
-A.5 Prompts
-The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A5 and A6.
+A.2 Examining Implicit shame
+We prompted GPT-4 chat to identify (a) if a given discourse expresses shame, and (b) identify the behavior or action behind shame, for a randomly sampled set of 100 samples without explicit mentions of shame/pride each from Hollywood and Bollywood.
+We asked two human annotators to evaluate the correctness of GPT-4 chat-generated outputs.
+A.3 Prompts
+The prompts designed to extract the subject experiencing self-conscious emotions, their gender, and the reason behind the emotions are provided in Tables A2 and A3.
 Prompts for norm discovery in Bollywood movie dialogues
@@
 Prompts for norm discovery in Hollywood movie dialogues
-"You are an experienced social scientist and you study Western society.
+"You are an experienced social scientist and you study American society.
 Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1.
@@
 Output: father, male, for getting the promotion"
-"You are an experienced social scientist and you study western societies.
+"You are an experienced social scientist and you study American societies.
 Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1.
@@
 Table A3: Hollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
-A.6 LIWC Correlation Results
+A.3.1 Annotation
+The annotation guidelines to verify the gender predicted by GPT-4 and the correctness of the reason are provided in Table A4.
+The annotators for Bollywood set and Hollywood set were Indian and American respectively.
+Both annotators were female, proficient in English language, and well-versed with social norms.
+During annotation, if the gender or the reason is unclear, the annotators were asked to label "not explicit".
+We only considered the cases where the gender was predicted to be either male or female.
+The task is objective and inter-annotator agreement was not computed.
+The annotators volunteered for the task and were not provided monetary compensation.
+Guidelines for Manual Evaluation
+Step-1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
+Step-2.
+Identify the gender.
+Check gender markers such as Mr/Mrs., s/he, him/her, etc.
+If the name is provided in the conversation, check if the name is likely to be a male name or female.
+If not clear, mark "not explicit".
+Step-3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
+Table A4: Guidelines for Annotation
+Gender Evaluation
+Incorrect Ambiguous
+Bollywood 8 15 (3 Female)
+Hollywood 5 10 (3 Female)
+Social Norms/Reason Evaluation
+Incorrect Ambiguous
+Bollywood 11 9
+Hollywood 2 1
+Table A5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
+Bollywood
+mastering a trick
+fulfilling father’s dreams
+provide care for old parents
+fiancee’s physical appearance
+his wealth
+Hollywood
+for being a hard worker regardless of the task
+being a brilliant student
+winning Olympic gold
+achievements and growth
+coming out as queer
+Table A6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
+A.4 LIWC Correlation Results
 Tables A7, A8, A9 and A10 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
@@
 The categories are arranged in decreasing order of correlation.
-A.7 Clustering Results
+A.5 Clustering Results
 Tables A13 and A14 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution.
@@
 Promiscuity being promiscuous, Having a love affair, shameless behavior and expressing love
-Inappropriate behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
+Inappropriate social behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
 Disobedience not upholding cultural values, disobedience or lack of respect, being disloyal
@@
 Gender Roles not behaving as per her husband’s expectations, behaving inappropriately in front of her daughter, going against her husband
-Sexual Harrassment assaulting a girl, physical assault on women, sexual assault
+Sexual Harassment assaulting a girl, physical assault on women, sexual assault
 Cowardice lack of courage, cowardice and inability to stand up for oneself, lack of pride and integrity
-Alcholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
+Alcoholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
 Non-conformity not conforming to gender norms, breaking gender norms, behaving inappropriately according to societal norms
@@
 Family norms Disrespecting family, causing harm and shame to family, causing difficulty and shame to family
-Table A11: Themes and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood.
+Lack of shame perceived shamelessness, being perceived as shameless, feeling embarrassed and ashamed
+Table A11: Themes ($n=26$) and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood.
+The cluster Lack of shame was removed since the reason behind shame was not evident.
+Clusters Lack of accountability and Inappropriate social behavior were merged together (and named ‘‘Lack of accountability’’) due to overlapping normative expectations.
 Theme Examples from cluster
@@
 Duty fulfilling duty, Accomplishing something great, saving the world
-Bravery & Nation bravery and sacrifice, bravery and service to nation, bravery and selflessness
+Bravery bravery and sacrifice, bravery and service to nation, bravery and selflessness
+Nation saving the country’s pride, making country proud, Making country proud
 Doing the "right" thing for doing the right thing, for being heroes, for having courage
-Son’s success son’s success and progress, son’s hard work and achievement, son’s determination
+Son’s achievements son’s success and progress, son’s hard work and achievement, son’s determination
 Family Honor maintaining dignity and pride, taking care of family pride, saving family’s pride
 Justice standing up for justice, fighting for justice, bringing justice
-Marriage-related marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
-Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
+Wedding marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
+Physical Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
 Resilience resilience and determination, overcoming struggles, enduring hardships without complaint
+Winning Sports achievement, working hard and achieving a spectacular result, winning something
 Ethnolinguistic Identity being Indian, defending the pride of Rajputs, being a Maharashtrian
-Table A12: Themes and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
+Table A12: Themes ($n=15$) and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
 Theme (Shame-related norms) Total Samples Hollywood Bollywood
-Accountability 802 185 617
+Lack of Accountability 802 185 617
 Accusation 70 0 70
@@
 Promiscuity 595 89 506
-Sexual behavior 407 33 374
+Inappropriate Sexual behavior 407 33 374
 Sexual Harassment 232 24 208
@@
 Finally, 24 clusters were considered.
-The total is slightly more than numbers in Table A3 since a norm could be mapped to multiple clusters.
+The total is slightly more than numbers in Table 5 since a reason could be mapped to multiple clusters.
 Theme (Pride-related norms) Total Samples Hollywood Bollywood
@@
 A total of fifteen clusters were generated with distance=5.
-Duplicates were removed for clustering.
-’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’.
-Finally, 13 clusters were considered for analysis.
-(a) Shame-related norms in Bollywood
-(b) Shame-related norms in Hollywood
-Figure A1: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood.
-$\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq.
+The total is slightly more than the numbers in Table 5 since a reason could be mapped to multiple clusters
+(a) Pride-related norms in Bollywood
+(b) Pride-related norms in Hollywood
+Figure A1: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood.
+$\Delta(Female-Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq.
 1 but for dimension gender).
-* indicates insignificant difference i.e., $p>0.05$
-(a) Shame-related norms in Bollywood
-(b) Shame-related norms in Hollywood
-Figure A2: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood.
-$\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq.
-1 but for dimension gender).
-* indicates insignificant difference i.e., $p>0.05$
+* indicates significant difference i.e., $p>0.05$
+Theme Statistic p-value
+Inappropriate Sexual behavior 5.0380 0.00001
+Gender roles 3.8262 0.0006
+Betrayal 3.2044 0.001
+Illegal activities 2.8258 0.005
+Disrespect 2.8284 0.005
+Sexual Harassment 2.8358 0.005
+Promiscuity 1.5553 0.1
+Accusation 3.8411 0.0006
+Stealing 1.6798 0.09
+Parent-related 1.8961 0.06
+Family norms 2.0813 0.04
+Non-conformity 1.4724 0.1
+Marriage-related 1.1817 0.2
+Privacy-related 0.9518 0.3
+Alcoholism -0.2920 0.8
+Immodesty -0.6634 0.5
+Disobedience -2.1438 0.03
+Lying/Deception -2.6508 0.008
+Cowardice -2.7746 0.006
+Incompetence -3.5187 0.002
+Social Etiquette -3.3461 0.002
+Harm -4.5140 0.00009
+Poverty -4.9028 0.00002
+Lack of accountability -4.6312 0.00007
+Table A15: Barnard Exact test with Yates correction for testing statistically significant occurrence of shame-related themes in Bollywood and Hollywood.
+Theme Statistic p-value
+Ethnolinguistic Identity 9.6185 0.0000
+Son’s Achievements 5.8025 0.0000
+Bravery 5.0368 0.0000
+Family Honor 4.7046 0.0000
+Nation 5.1183 0.0000
+Wedding 5.3898 0.0000
+Resilience 1.4959 0.1370
+Justice 0.1441 0.8963
+Family Roles -0.2560 0.8024
+Physical Appearance -0.6420 0.5223
+Winning -2.8913 0.0038
+Achievement -3.5425 0.0004
+Self-identity -4.9482 0.0000
+Doing the "right" thing -10.4936 0.0000
+Duty -8.5488 0.0000
+Table A16: Barnard Exact test with Yates correction for testing statistically significant occurrence of pride-related themes in Bollywood and Hollywood.

### v5

@@
 We examine how and why shame and pride are expressed across cultures using a blend of psychology-informed language analysis combined with large language models.
-We find significant cross-cultural differences in shame and pride expression aligning with known cultural tendencies of the USA and India -- e.g., in Hollywood, shame-expressions predominantly discuss self whereas Bollywood discusses shame toward others.
-Pride in Hollywood is individualistic with more self-referential singular pronouns such as I and my whereas in Bollywood, pride is collective with higher use of self-referential plural pronouns such as we and our.
-Lastly, women are more sanctioned across cultures and for violating similar social expectations e.g.
-promiscuity.
+We find significant cross-cultural differences in shame and pride expression aligning with known cultural tendencies of the USA and India – e.g., in Hollywood, shame-expressions predominantly discuss self whereas shame is expressed toward others in Bollywood.
+Women are more sanctioned across cultures and for violating similar social expectations.
 Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice
+Sunny Rai, Khushang Jilesh Zaveri, Shreya Havaldar, Soumna Nema, Lyle H.
+Ungar & Sharath Chandra Guntuku University of Pennsylvania, Johns Hopkins University, Mahindra University {sunnyrai, sharathg}@seas.upenn.edu
 1 Introduction
@@
 (2011) and self-esteem Pyszczynski et al.
-(2004) reliably from language -- enabling effective human-computer interactions in domains such as AI-driven psychotherapy Resendiz and Klinger (2023); Demszky et al.
+(2004) reliably from language – enabling effective human-computer interactions in domains such as AI-driven psychotherapy Resendiz and Klinger (2023); Demszky et al.
 (2020).
-Examining behaviors evoking shame and pride will reveal culture-specific societal expectations and attitudes (See Table 1) -- paramount to building culturally competent AI models for diverse users Talat et al.
+Examining behaviors evoking shame and pride will reveal culture-specific societal expectations and attitudes (See Table 1) – paramount to building culturally competent AI models for diverse users Talat et al.
 (2021); Atari et al.
 (2023).
-Look, Alok is in love with me and I love him too.
-Shameless one! Have you lost all your shame?
-What is there to be ashamed in this? Don’t sisters love their brothers in this vicinity?
-$\Rightarrow$ Incestuous relationship evoke shame.
+And should we bow before others begging....them to marry our daughters?
+This shall not happen.
+Neither will the girls be alive here nor shall....we be ashamed of ourselves.
+You cannot kill the life which God has given.
+I won’t let you commit the sin.
+$\Rightarrow$ Not able to marry off their daughters evokes shame.
 Sister-in-law! Congrats, sister-in-law! Big brother has started working!
@@
 Table 1: Excerpts of dialogues expressing explicit social emotions shame and pride indicating actions and social approval.
-For our study, we select two top-grossing movie industries, i.e., Hollywood and Bollywood  [^1] https://nairametrics.com/2024/08/14/top10-fastest-growing-film-industries-in-the-world/.
+For our study, we select two top movie industries, i.e., Hollywood and Bollywood [^1] https://en.wikipedia.org/wiki/Film_industry.
 Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988).
@@
 individualism) thus presents a rich ground for understanding variations in their beliefs and values.
-Our approach blends psychology-informed language analysis with state-of-the-art Large Language Model (LLM) to delineate interpretable psychosocial constructs such as self-focus, morality, etc.
-in shame/pride-related expressions and extract implicit social expectations behind them.
-This paper makes the following research contributions: (a) We build a cross-cultural dataset of over 10k shame/pride-related dialogues with underlying social expectations  [^2] Link will be provided upon acceptance, (b) We demonstrate significant cross-cultural variations in how and why shame and pride are expressed in movies, and (c) We illustrate that women are more subjected to social sanctions than men across cultures and for violating similar social expectations.
-To the best of our knowledge, this is the first empirical study analyzing the psycho-social constructs underlying shame/pride-related expressions across cultures and the social expectations behind them.
+Our approach blends psychology-informed language analysis with state-of-the-art Large Language Model (LLM) to delineate interpretable psychosocial constructs such as self-focus and morality in shame/pride-related expressions and extract implicit social expectations behind them.
+This paper makes the following research contributions: (a) We develop a cross-cultural dataset of over 10k shame/pride-related dialogues with underlying social expectations [^2] https://github.com/Khushangz/Cross-Cultural-Social-Norms-Dataset/, (b) We demonstrate cross-cultural variations in how and why shame and pride are expressed in movies, and (c) We illustrate that women are more subjected to social sanctions than men across cultures and for violating similar social expectations.
+To the best of our knowledge, this is the first empirical study analyzing the psychosocial constructs underlying shame/pride-related expressions across cultures and the social expectations behind them.
 2 Movies Subtitles Corpora
@@
 The natural conversation style between characters in movies can reveal the social power dynamics (e.g., boss-employee, father-daughter) and gender roles.
-Another potential alternative is social media posts however, they are (a) skewed to social situations specific to young demographic groups from Western countries, and (b) are second-hand reports that do not reveal fully the social dynamics of how shame/pride are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
+Another potential alternative is social media posts however, they are (a) skewed to social situations specific to young demographic groups from Western countries, and (b) are second-hand reports that do not reveal fully the social dynamics of how shame/pride are expressed in natural conversations, for instance, between a boss and an employee.
 We collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table 2 for data distribution).
@@
 Bollywood movie subtitles were collected from www.Bollynook.com and Hollywood movie subtitles were from the publicly available Kaggle movie subtitles dataset [^3] https://www.kaggle.com/datasets/adiamaan/movie-subtitle-dataset.
-The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries.
 The year mapping was performed to ensure a similar period for collected movies.
+The year of release for movies was verified by either parsing subtitle file names having a release year or probing Wikipedia entries.
 Hollywood Bollywood
@@
 Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud).
-Control is the group of dialogues without words shame, pride, and their derivatives.
+Control comprises dialogues without words shame, pride, or their derivatives.
 2.1 Extracting Shame and Pride related Expressions
@@
 We adopted a keyword search-based approach (See Appendix A1) to identify dialogues along with the previous and the next two lines for situational context (See Table 1).
-In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context.
-One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues with phrases such as "what a shame", "it’s a shame", "proudly presents".
-Detecting implicit shame/pride-related expressions in textual discourse is an innately complex task.
-Computational systems for detecting social emotions such as shame and guilt underperform significantly compared to basic emotions such as joy Demszky et al.
-(2020); Resendiz and Klinger (2023) due to internal self-evaluation associated with social emotions.
-We tested the GPT-4 language model Achiam et al.
-(2023) to detect implicit expressions of shame in given movie dialogues without explicit mentions of shame and pride.
+For short dialogues such as monosyllabic responses in spoken conversations, we appended an extra previous and next line to context.
+One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues with conventional phrases such as "what a shame", "it’s a shame", "proudly presents".
+Detecting implicit shame/pride-related expressions in textual discourse is a complex task.
+Computational systems for detecting social emotions such as shame and guilt underperform significantly compared to basic emotions such as joy due to internal self-evaluation associated with social emotions Demszky et al.
+(2020); Resendiz and Klinger (2023).
+We used GPT-4 language model Achiam et al.
+(2023) to detect implicit expressions of shame in given movie dialogues.
 We recruited two human annotators to determine the agreement with GPT-4 generated labels (See Appendix A.2 for details).
-On a randomly sampled 100 situations each from H/Bollywood, the model returned 20 samples to be expressing shame in Bollywood and 5 samples in Hollywood.
-We found poor agreement between human annotators (4 out of 20 samples in Bollywood and 1 out of 5 samples in Hollywood where both annotators agreed with GPT-4).
+On a randomly sampled 100 situations each from H/Bollywood, the model labeled 20 samples to be expressing shame in Bollywood and 5 samples in Hollywood.
+However, the agreement between human annotators was low (4 out of 20 samples in Bollywood and 1 out of 5 samples in Hollywood where both annotators agreed with GPT-4).
 Human annotators across cultures tend to cross-label shame and guilt Troiano et al.
@@
 One such example from Hollywood is provided below:
-‘‘about feeding the poor, I’ve never done any of that. God tells us to love everybody. I’ve hated people… my family, my family, ’’
+about feeding the poor, I’ve never done any of that. God tells us to love everybody. I’ve hated people… my family, my family,
 Here, the speaker is probably expressing guilt, shame, or both.
 Given subjective beliefs and poor agreement with generated outputs, we adopted a high precision-low recall strategy and limited our further analysis to explicit mentions of shame and pride.
-For our analysis, we created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood.
+We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood.
 Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table 2 for dialogue distribution).
 Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
-Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
-Moral (wrong, excuse, decent, honest, duty) 0.110 $2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.161
-$1^{st}$ person sing. pronouns (i, me, i’m, my, i’ll) 0.102 Social References (you, your, he, her, him ) 0.148
-Sadness (crying, cry, sob, lonely, sad) 0.096 Moral (wrong, innocent, duty, decent, excuse) 0.102
-$2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.080 Female ( her, she, girl, she’s, mom) 0.060
-Past focus (was, did, were, been, didn’t ) 0.069 Family (son, married, uncle, dad, mom) 0.055
-All or none (no, all, nothing, never, yes) 0.055 Sexual (chaste, lust, sex, sexy, pimp) 0.051
-Family (son, dad, baby, mom, mama) 0.054 Conflict (kill, killed, accusing, killing, cruel) 0.024
-Social References (you, your, we, he, you’re) 0.050 Swear words (hell, bloody, idiot, damn, ass) 0.024
-Anxiety (fear, afraid, worry, terrified, scared) 0.042 Anger (mad, angry, hate, cruel, argue) 0.022
-Illness (sick, pain, pains, flu, sickly) 0.038 Present focus (is, are, don’t, i’m, aren’t) 0.021
-Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values).
-Note, $1^{st}$ person sing pronouns are strongly correlated with Hollywood-shame whereas the correlation with $2^{nd}$ person pronouns and social references are up to 3 times stronger in Bollywood compared to Hollywood.
-See Table A7 and A9 for the complete set of correlations.
 3 Approach
 Fig.
-1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting LLM approach to extract reasons behind them and how they differ between India and the U.S.A.
-3.1 Vocabulary Approach: LIWC
+1 illustrates our study design comprising two key approaches (a) using a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting LLM approach to extract reasons behind them and how they differ between India and the U.S.A.
+3.1 Vocabulary Approach
 Linguistic Inquiry of Word Count (LIWC) Boyd et al.
-(2022) is a corpus analysis tool widely used in psychology to identify psychological categories of words such as "self-focus" (e.g., I, my, mine) and "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
-To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psycho-social categories from the dialogues and examined their correlation with shame and pride compared to the control set.
+(2022) is a corpus analysis tool to identify psychological constructs such as "self-focus" and "morality", and therefore, is widely used for examining social behaviors such as self-regulation and conformity.
+To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psychosocial categories in LIWC from the dialogues and examined their correlation with shame and pride compared to the control set.
 The search keywords used for building social emotions corpus (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
-3.2 Pretrained LLM: GPT-4 Chat
+3.2 Prompting Approach
 Identifying why shame/pride is expressed in discourse requires multicultural world knowledge.
@@
 Unlike movie dialogues, reasons extracted from GPT-4 chat are short phrases (See Table 4) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture-specific language is needed.
-Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
-Social References (you, your, we, he, you’re) 0.131 Drive (we, our, us, sir, work) 0.122
-Family (son, dad, baby, mom, mama) 0.121 Achieve (work, better, win, best, try) 0.098
-$1^{st}$ person sing. pronouns (i, i’m, me, my, i’ll) 0.095 Power (sir, own, respect, kill, power) 0.091
-$2^{nd}$ person pronouns (you, your, you’re, you’ve, yourself) 0.091 Social References (you, your, he, we, our) 0.085
-Male References (he, his, him, man, son) 0.084 Moral (wrong, duty, brave, arrogant, useless) 0.075
-Drive (we, our, us, we’re, dad) 0.076 Reward (win, won, glory, success, successful) 0.071
-Ethnicity (American, Irish, Chinese, German, Christian) 0.052 Political (nation, army, sultan, president, dynasty) 0.064
-Reward (win, won, winner, successful, earned) 0.048 $1^{st}$ person pl. pronouns (we, our, us, we’ll, let’s) 0.060
-Achieve (work, better, best, trying, try) 0.045 Family (son, papa, married, dad, uncle) 0.059
-Power (sir, own, war, strong, mighty) 0.044 Ethnicity (Indian, Indians, British, Hindi, caste) 0.054
-Table 4: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values).
-Note the contrast, Achievement-related pride in Bollywood, Self-centered pride vs We-centered pride, Family in Hollywood vs Family in Bollywood See Tables A8 and A10 for all significantly correlated categories.
+Figure 2: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values).
+Note, $1^{st}$ person sing pronouns are strongly correlated with Hollywood-shame whereas the correlation with $2^{nd}$ person pronoun and social references are up to 3 times stronger in Bollywood compared to Hollywood.
+See Table A7 and A9 for the complete set of correlations.
+Figure 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values).
+Note the contrast, Achievement-related & We-centered pride in Bollywood vs Social & Self-centered pride in Hollywood.
+See Table A8 and A10 for the complete set of correlations.
+4 Results
+4.1 How is shame/pride expressed across cultures?
+Shame is associated with negative emotions, power, and morality in both movie industries (See Tables A7 and A9) however, significant cultural nuances exist in its manifestation (See Fig.
+2).
+In Bollywood, shame-related expressions are other-oriented as indicated by 2nd person pronouns and social references.
+Female and sexual categories are correlated with shame exclusively in Bollywood potentially indicating the honor system in collectivist societies Caffaro et al.
+(2014).
+Other psychosocial categories such as swearing and conflict with present focus reveal the role of shame in enforcing conformity.
+In Hollywood, shame expressions are self-focused with remorse (e.g., sadness, anxiety) and have past-focused language.
+Illness is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
+A similar dichotomy is observed for pride (See Fig.
+3).
+In Bollywood, pride is achievement/power focused and "We"-centered.
+Pride is associated exclusively with men in Bollywood.
+Other psychosocial categories such as moral and politics are more strongly associated with pride.
+In Hollywood, pride is more self-centered and expressed for family more.
+Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood.
+Pride has a positive undertone in Hollywood whereas in Bollywood, we speculate that it relates to honor (e.g., protecting family’s pride, bringing pride to family) and does not have a positive undertone.
+4.2 How do social expectations behind shame/pride vary across cultures?
+Cross-Cultural Shame/Pride Dataset
+For the Bollywood set, GPT-4 predicted reasons (also implicit social expectations) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999.
+For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805.
+Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as children.
+As a result, a lower number of reasons is associated with pride compared to shame.
+GPT-4 chat predicted a total of 10,445 social expectations (See Table 3 for distribution).
+Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 4 for shame-related expectations and A6 for pride-related expectations).
+4.2.1 Manual Evaluation
+Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" for a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood.
+An Indian annotator aware of social roles and expectations in Indian society labeled the Bollywood set.
+Likewise, an American annotator labeled the Hollywood set.
+For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table A5 for more details).
+There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
 Bollywood Hollywood
@@
 The duplicate reasons are not removed as their frequency reflects their prevalence and is useful for estimating gender association.
-4 Results
-4.1 How is shame/pride expressed across cultures?
-Shame is associated with negative emotions, power, and morality in both movie industries (See Table 3), however, significant cultural nuances exist in its manifestation.
-In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language.
-In Bollywood, psychosocial categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict).
-Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al.
-(2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
-A similar dichotomy is observed for pride (See Table 4).
-Pride in Hollywood is more centered toward family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Table A8).
-Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood.
-Pride is exclusively positive (see POSEMO in Table A10) in Hollywood whereas in Bollywood, we suppose that it relates to honor more (e.g., protecting family’s pride, bringing pride to family) and is not always used in a positive context.
-4.2 How do social expectations behind shame/pride vary across cultures?
-Cross-Cultural Shame/Pride Dataset
-For the Bollywood set, GPT-4 chat predicted reasons (also implicit social expectations) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999.
-For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805.
-Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of reasons is associated with pride compared to shame.
-GPT-4 chat predicted a total of 10,445 social expectations (See Table 5 for distribution).
-Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 6 for shame-related expectations and A6 for pride-related expectations).
-4.2.1 Manual Evaluation
-Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood.
-The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society.
-Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social expectations of the U.S.A.
+4.2.2 Thematic Analysis
+Twenty-four clusters for shame-related reasons and fifteen for pride-related reasons were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters).
+The clusters were manually assigned a theme (as depicted on the Y-axis of Fig.
+4 and 5) after analyzing the ten closest samples based on cosine distance from the centroid of the cluster (See Table A11 and A12 for examples in each cluster).
+We computed the relative association for each cluster with Bollywood and Hollywood using eq.
+1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
 Bollywood
@@
 Table 4: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
-For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table A5 for more details).
-There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
-4.2.2 Thematic Analysis
-Twenty-four clusters for shame-related reasons and fifteen for pride-related reaosns were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters).
-The clusters were manually assigned a theme (as depicted on the Y-axis of Fig.
-2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples in each cluster).
-We computed the relative association for each cluster with Bollywood and Hollywood using eq.
-1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
 $\vec{\Delta}=\forall_{t_{i}\in themes}\frac{\mathcal{D}_{bolly_{t_{i}}}}{{\mathcal{D}_{bolly}}}-\frac{\mathcal{D}_{holly_{t_{i}}}}{\mathcal{D}_{holly}}$ (1)
-Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 2 and 3).
+Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 4 and 5).
 Figure 4: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms.
-* indicates significant difference i.e., $p>0.05$.
+* indicates significant difference i.e., $p<0.05$.
 See Table A15 for statistics and p-value.
@@
 Themes such as lack of accountability and poverty are more common in Hollywood whereas inappropriate sexual behavior and gender roles are more prevalent in Bollywood (See Fig.
-2).
+4).
 Unsurprisingly, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, causing harm and incompetence evoke shame in Hollywood.
@@
 Figure 5: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
-* indicates significant difference i.e., $p>0.05$.
+* indicates significant difference i.e., $p<0.05$.
 See Table A16 for statistics and p-value.
-Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
+Figure 6: Relative association ($\Delta$) of social emotions attributed to male and female. A higher positive score indicates a stronger association of gender with pride.
 $\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
-Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related words where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender.
+Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related words where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents the number of movie dialogues spoken by $g$.
 As depicted in Fig.
-4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented.
+6, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented.
 Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
 Social expectations evoking shame are similar across movie industries (See Fig.
-5 and 6).
-Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries.
+7 and 8).
+Sexuality (e.g., promiscuity, immodesty) is the dominant theme for sanctioning women in both movie industries whereas males are shamed for incompetency.
 Women across movie industries express more pride in family roles.
@@
 A1 for gender differences in pride).
-Figure 5: Gender differences in shame-related themes in Bollywood movies. * indicates significant difference i.e., $p>0.05$.
-Figure 6: Gender differences in shame-related themes in Hollywood movies. * indicates significant difference i.e., $p>0.05$.
+Figure 7: Gender differences in shame-related themes in Bollywood movies. * indicates significant difference i.e., $p<0.05$.
+Figure 8: Gender differences in shame-related themes in Hollywood movies. * indicates significant difference i.e., $p<0.05$.
 5 Discussion
-We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related expressions with implicit reasons behind them.
+We release a cross-cultural dataset of shame- and pride-related expressions with implicit reasons behind them.
 To the best of our knowledge, the only related corpus is the GoEmotions dataset Demszky et al.
 (2020) comprising 817 samples for embarrassment and 452 samples for pride.
-This corpus can serve the task of developing computational models for detecting shame/pride expressions and recognizing how appraisals of such emotions vary -- enabling culturally cognisant human-computer conversations for diverse users Kim et al.
+Our dataset can help develop computational models for detecting shame/pride expressions and recognizing how appraisals of such emotions vary – enabling culturally cognisant human-computer conversations for diverse users Kim et al.
 (2011) and cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
-Shame is a highly undesirable self-focused emotion emphasizing incompetency and failures and is rarely used in the U.S.
-Cohen (2003); Boiger et al.
-(2013) whereas Eastern cultures use shame more frequently and have a desirable affect Lim (2016).
-Notably, we also observe infrequent use of shame in Hollywood movies and self-focus (reflecting internal shame) (See Table 2 and 3 for frequency distribution and psychosocial constructs).
-Unlike in Western nations, shame is interdependent in collectivist communities Wong and Tsai (2007).
-The empirical analyses reveal other-focused (also, the public nature of shame) in Bollywood movies (See Table A7).
-Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs.
-present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007).
-Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial).
-It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10 for all significant categories correlated with pride).
+Shame is a highly undesirable emotion emphasizing incompetency and failures in the U.S.
+and used sparingly Cohen (2003); Boiger et al.
+(2013) whereas Eastern cultures use shame more frequently and it is known to have a desirable affect Lim (2016).
+Notably, we also observe that shame is 4.5 times more common in Bollywood compared to Hollywood.
+We also observe that shame is self-focused (reflecting internal shame) in Hollywood (See Table 2 and Fig.
+2 for frequency distribution and psychosocial constructs).
+In contrast, shame is interdependent in collectivist communities Wong and Tsai (2007).
+The empirical analyses reveal other-focused (also, the public nature of shame) in Bollywood movies (See Fig.
+3).
+Moreover, the contrasting tenses coupled with varying emotions, i.e., past-focus + sadness in Hollywood vs.
+present-focus + conflict in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007).
+Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings that the significance of "success" grows as a society becomes more individualistic Cohen (2003).
+In Bollywood, pride is collective and male-focused (we, achievement, male-focused) Khadilkar et al.
+(2022).
+It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicate more materialistic pride, compared to personal pronouns and social references in Hollywood which reflect more personal pride.
 Eliciting reasons leading to shame and pride expressions revealed social expectations across cultures.
@@
 returning calls after date night in Hollywood).
-To tease apart ways in which cultures differ, we first considered mapping social expectations/generated reasons to Schwartz’s Theory of Values Schwartz (2012).
+To tease apart cultural differences, we first considered mapping social expectations/generated reasons to Schwartz’s Theory of Values Schwartz (2012).
 However, social expectations were found to have contrasting values depending on the culture.
 Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context.
-To effectively understand the cross-cultural variations in social expectations underlying shame/pride, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters.
+We thus performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters.
 The situations connected to shame and pride in the U.S.
-society (i.e., lacking accountability, poverty, harm) vs Indian society (e.g., inappropriate sexual behavior, gender roles, betrayal - see Fig 2 and 3) align with known cultural tendencies (i.e., individualist vs collectivist) of both nations Triandis (1988).
+society (i.e., lacking accountability, poverty, harm) vs Indian society (e.g., inappropriate sexual behavior, gender roles, betrayal - see Fig 4 and 5) align with known cultural tendencies (i.e., individualist vs collectivist) of both nations Triandis (1988).
 The vision of safe and accountable AI is centered on LLMs’ moral and value alignment.
-The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al.
+The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or identifying underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al.
 (2021); Fung et al.
@@
 (2023).
-However, most of the social situations and the human annotators employed to label those situations reflect English beliefs and ethics.
-Our approach to extracting culture-specific social expectations and attitudes using social emotions overcomes this limitation.
-While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are similar to patterns seen during LIWC analysis, validating our approach for overcoming Anglocentric bias in LLMs for norm discovery Ziems et al.
-(2022).
-Lastly, we examined how social expectations associated with shame/pride are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021).
-We demonstrated that shame, a negative self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, endorsing increased social value, is used for more men than women in movies.
+However, a majority of social situations and the human annotators employed to label those situations reflect English beliefs and ethics.
+Our approach for extracting culture-specific social expectations and attitudes using social emotions overcomes this limitation.
+While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are similar to patterns seen during LIWC analysis, validating our approach for overcoming Anglocentric bias in LLMs for norm discovery.
+Lastly, we examined how social expectations associated with shame/pride are attributed to men and women across cultures influencing their social behaviors.
+For example, women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021).
+We demonstrated that shame, a negative self-conscious emotion expressing devaluation, is targeted toward women more than men, and pride, a social emotion endorsing social value, is more used for men than women.
 The similarity in social expectations (e.g.
-regulating sexual behaviors in women vs shaming incompetency in men) across cultures is surprising (See Fig 5 and 6).
-It is important to characterize the source and acquired social expectations before their use for achieving alignment in LLMs otherwise we risk social bias.
+regulating sexual behaviors of women vs shaming incompetency in men) across cultures is surprising.
+We thus feel that it is important to characterize social biases in the training data before their use for aligning LLMs.
 6 Conclusion
-We introduce a multi-cultural dataset of shame/pride-related expressions and the underlying social expectations.
-This study demonstrates (a) cross-cultural linguistic differences in shame and pride-related expression, offering insights into their functions across cultures, (b) cultural dichotomy in social expectations, and (c) more "social sanctions" and fewer "endorsements" to the female gender in movies.
+We introduced a multi-cultural dataset of shame/pride-related expressions and the underlying social expectations.
+This study demonstrates (a) cross-cultural linguistic differences in shame and pride-related expression, offering insights into their functions across cultures, (b) cultural dichotomy in social expectations, and (c) more "social sanctions" and fewer "endorsements" to women in movies.
 Future work can utilize our dataset for culturally aligned LLMs and build social emotion perception and appraisal in human-computer interactions.
@@
 Social Impact and Ethics Statement
-Social norms discovery is a crucial component in programs [^6] ALIGN-https://www.alignplatform.org/learning-collaborative designed for social and behavioral change to promote equity, social justice, and well-being Mauduy et al.
+Social norms discovery is a crucial component of social and behavioral change programs [^6] ALIGN-https://www.alignplatform.org/learning-collaborative promoting equity, social justice, and well-being Mauduy et al.
 (2022); Bonan et al.
@@
 Moreover, countries like India and America contain a mix of cultures.
-The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North India vs South India).
+The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North vs South in India).
 Movies also exaggeratedly depict the world around us (e.g., weddings, criminal activities, sexual abuse, etc.), and we caution against stereotyping cultures based on movie-based norms.
 The dominant language in Bollywood movies is Hindi and our analysis is based on their English translations which may not always be accurate, especially when the discourse is about concepts native to a culture.
-Relatedly, the LIWC may not have high coverage for such concepts.
+LIWC dictionary may also lack complete coverage for such concepts.
 We did not compare the movie genre and acknowledge that situational/unrealistic norms (e.g., science fiction, comedy, etc.) could exist.
 Regardless, social norms associated with shame and pride are still relatable and reflect the target audience’s beliefs.
-It is also worth noting that gender norms in movies may not fully reflect the norms of today’s society.
-This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
+This study is conducted over a thirty-year long period and the aggregated norms may not truly reflect the current trends.
 We acknowledge that the excerpts expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America).
@@
 Acknowledgements
-We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in refining this manuscript.
+This work was supported in part by the Penn Global Engagement Research Fund. We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance with writing.
 Appendix A Appendix
@@
 We asked two human annotators to evaluate the correctness of GPT-4 chat-generated outputs.
+See Table A4 for annotation guidelines.
 A.3 Prompts
-The prompts designed to extract the subject experiencing self-conscious emotions, their gender, and the reason behind the emotions are provided in Tables A2 and A3.
+The prompts designed to extract the subject experiencing shame and pride, their gender, and the reason behind the emotions are provided in Tables A2 and A3.
 Prompts for norm discovery in Bollywood movie dialogues
@@
 2.
-Identify the primary reason for shaming The answer should be short and in CSV format.
-$<$ shamed, gender, reason$>$ Input: And should we bow before others begging....them to marry our daughters?
+Identify the primary reason for shaming.
+If not explicit, return NA.
+The answer should be short and in CSV format.
+$<$ who_shame, gender, reason$>$ Input: And should we bow before others begging….them to marry our daughters?
 This shall not happen.
-Neither will the girls be alive here nor shall....we be ashamed of ourselves.
+Neither will the girls be alive here nor shall….we be ashamed of ourselves.
 You cannot kill the life which God has given.
@@
 And the thieves too Politics is in a great mess Shame on this system.
-There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’ Output: System, NA, poor law and regulations"
+There’s no democracy Get rid of these politicians The gong has struck..”Our hearts are swaying to it’s beats” Output: System, NA, poor law and regulations"
 Table A2: Bollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
@@
 Output: father, male, for getting the promotion"
-"You are an experienced social scientist and you study American societies.
+"You are an experienced social scientist and you study American society.
 Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1.
@@
 Identify the primary reason for shaming.
+If not explicit, return NA.
 The answer should be short and in CSV format.
-$<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100.
+$<$ who_shame, gender, reason$>$ Input: You still owe me 100.
 Remember?
@@
 A.4 LIWC Correlation Results
-Tables A7, A8, A9 and A10 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
+Tables A7, A8, A9 and A10 contain positively correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, Pearson r and confidence interval.
 LIWC Categories Top-5 words r p-value 95% CI
@@
 EMOTION (love, good, bad, happy, crazy) 0.168 0.000 [0.152, 0.185]
-YOU (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
+$2^{nd}$ person pronouns (YOU) (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
 Social references (SOCREFS) (you, your, he, her, him) 0.148 0.000 [0.132, 0.165]
@@
 PRONOUN (you, i, me, your, my) 0.040 0.000 [0.023, 0.057]
-AUXVERB (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
+Auxiliary verbs (AUXVERB) (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
 CONFLICT (kill, killed, accusing, killing, cruel) 0.024 0.007 [0.007, 0.041]
@@
 FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 0.018 [0.005, 0.038]
-Table A7: Psychosocial categories significantly correlated (p<0.05) with shame in Bollywood dialogues.
+Table A7: Psychosocial categories positively correlated ($p<0.05$) with shame in Bollywood.
 p-values were corrected using Benjamini-Hochberg correction.
@@
 FUNCTION (you, the, i, of, to) 0.032 0.001 [0.013,0.050]
-YOU (you, your, you’re, you’ve, you’ll) 0.030 0.003 [0.011,0.048]
+$2^{nd}$ person pronouns (YOU) (you, your, you’re, you’ve, you’ll) 0.030 0.003 [0.011,0.048]
 THEY (they, their, them, they’re, they’ll) 0.029 0.003 [0.011,0.048]
@@
 Positive Emotion (EMO_POS) (love, good, happy, happiness, smile) 0.020 0.049 [0.001,0.038]
-Table A8: Psychosocial categories significantly correlated (p<0.05) with pride in Bollywood dialogues.
+Table A8: Psychosocial categories positively correlated ($p<0.05$) with pride in Bollywood.
 p-values were corrected using Benjamini-Hochberg correction.
@@
 MORAL (wrong, excuse, decent, honest, duty) 0.110 0.000 [0.084, 0.136]
-I (i, me, i’m, my, i’ll) 0.102 0.000 [0.076, 0.128]
+$1^{st}$ person sing. pronouns (I) (i, me, i’m, my, i’ll) 0.102 0.000 [0.076, 0.128]
 Sadness (EMO_SAD) (crying, cry, sob, lonely, sad) 0.096 0.000 [0.070, 0.122]
@@
 COGNITION (no, not, all, know, but) 0.032 0.033 [0.005, 0.058]
-Table A9: Psychosocial categories significantly correlated (p<0.05) with shame in Hollywood dialogues.
+Table A9: Psychosocial categories positively correlated (p<0.05) with shame in Hollywood.
 p-values were corrected using Benjamini-Hochberg correction.
@@
 FUNCTION (you, i, the, of, to) 0.100 0.000 [0.077,0.123]
-I (i, i’m, me, my, i’ll) 0.095 0.000 [0.072,0.118]
+$1^{st}$ person sing. pronouns (I) (i, i’m, me, my, i’ll) 0.095 0.000 [0.072,0.118]
 YOU (you, your, you’re, you’ve, yourself) 0.091 0.000 [0.068,0.114]
@@
 WORK (work, job, school, deal, company) 0.028 0.033 [0.005,0.051]
-Table A10: Psychosocial categories significantly correlated (p<0.05) with pride in Hollywood dialogues with 95% confidence intervals.
+Table A10: Psychosocial categories positively correlated (p<0.05) with pride in Hollywood.
 p-values were corrected using Benjamini-Hochberg correction.
@@
 A.5 Clustering Results
-Tables A13 and A14 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution.
+Tables A13 and A14 contain the manually annotated Cluster Themes, the total number of samples in each cluster, and Bollywood vs Hollywood distribution.
 The distance was set to $5$ and the duplicates were removed.
@@
 The cluster Lack of shame was removed since the reason behind shame was not evident.
-Clusters Lack of accountability and Inappropriate social behavior were merged together (and named ‘‘Lack of accountability’’) due to overlapping normative expectations.
+Clusters Lack of accountability and Inappropriate social behavior were merged together (and named “Lack of accountability”) due to overlapping normative expectations.
 Theme Examples from cluster
@@
 1 but for dimension gender).
-* indicates significant difference i.e., $p>0.05$
+* indicates significant difference i.e., $p<0.05$
 Theme Statistic p-value